### v3

# Splittings of One-Ended Groups with One-Ended Halfspaces

Abstract
We introduce the notion of halfspaces associated to a group splitting, and investigate the relationship between the coarse geometry of the halfspaces and the coarse geometry of the group. Roughly speaking, the halfspaces of a group splitting are subgraphs of the Cayley graph obtained by pulling back the halfspaces of the Bass–Serre tree. Our first theorem shows that (under mild conditions) any splitting of a one-ended group can be upgraded to a splitting where all the halfspaces are one-ended. Our second theorem demonstrates that a one-ended group usually has a JSJ splitting where all the halfspaces are one-ended. And our third theorem states that if a one-ended finitely presented group $G$ admits a splitting such that some edge stabilizer has more than one end, but the halfspaces associated to the edge stabilizer are one-ended, then $H^{2}(G,\mathbb{Z}G)\neq\{0\}$; in particular $G$ is not simply connected at infinity and $G$ is not an $n$-dimensional duality group for $n\geq 3$.
1 Introduction
In 1982, B. Jackson [Jac82b] proved that amalgamated products and HNN-extensions of one-ended finitely presented groups over finitely generated groups with more than one end are not simply connected at infinity. When the base groups are not one-ended, little was known. Exotic one-ended groups can be obtained as amalgamated products (HNN-extensions) with infinite-ended base groups and infinite-ended edge groups. These groups may or may not be simply connected at infinity. In this paper, instead of considering splittings with one-ended vertex groups we consider splittings with one-ended halfspaces. Given a finitely generated group $G$ and a splitting $G\curvearrowright T$ (here $T$ denotes the Bass–Serre tree) with finitely generated edge stabilizers, halfspaces in $T$ induce halfspaces in the Cayley graph of $G$ (see Section 2 for precise definitions). Each halfspace is a connected subgraph of the Cayley graph, and its quasi-isometry type is independent of the choice of finite generating set for $G$ (Lemma 2.4). If $G$ is one-ended one may expect that the halfspaces are also one-ended, but this is not necessarily the case – see Example 7.2. However, our first theorem shows that under mild conditions one can upgrade a splitting of a one-ended group to one with one-ended halfspaces. This result is in analogy with results about one-ended halfspaces for cubulated groups that appear in [She23].
Theorem 1.1.
Let $G\curvearrowright T$ be a non-trivial splitting with $G$ one-ended and finitely generated. Suppose the edge stabilizers are finitely generated and accessible. Then there is a non-trivial splitting $G\curvearrowright T^{\prime}$ with minimal action such that:
  1. (1)
    The halfspaces of $G\curvearrowright T^{\prime}$ are one-ended.
  2. (2)
    Edge stabilizers (resp. vertex stabilizers) of $T^{\prime}$ are finitely generated and are subgroups of the edge stabilizers (resp. vertex stabilizers) of $T$.
  3. (3)
    For each edge $e^{\prime}$ in $T^{\prime}$ there exists an edge $e$ in $T$ such that $G_{e^{\prime}}$ is a vertex stabilizer in some finite splitting of $G_{e}$ over finite subgroups.
As a corollary of Theorem 4.1 (which is the main technical result behind Theorem 1.1) we also have the following result.
Corollary 1.2.
If a one-ended group splits non-trivially over a two-ended group then the halfspaces of this splitting are one-ended.
As an elementary example of Corollary 1.2 one can consider an essential simple closed curve in a closed hyperbolic surface $S$ and the associated splitting of $\pi_{1}(S)$ over $\mathbb{Z}$ with free vertex groups; the lift of the curve to the universal cover $\mathbb{H}^{2}$ splits $\mathbb{H}^{2}$ into two halfspaces each of which is a half-plane.
There is a long history in the literature of studying group splittings and of finding ways to upgrade a given splitting to one that has better properties or that carries more information about the group. Probably the most well known example of this is the notion of a JSJ splitting or JSJ tree, due to Rips–Sela [RS97] and Dunwoody–Sageev [DS99]. In its most general form, if we have a group $G$ and a family $\mathcal{A}$ of subgroups of $G$ that is closed under conjugating and taking subgroups, then there is a notion of JSJ tree of $G$ over $\mathcal{A}$ due to Guirardel–Levitt [GL17] (see Definition 4). Our next theorem shows that, under mild conditions on $\mathcal{A}$, any one-ended group has a JSJ splitting with one-ended halfspaces.
Theorem 1.3.
Let $G$ be a finitely generated one-ended group and let $\mathcal{A}$ be a family of subgroups of $G$ that is closed under conjugating and taking subgroups.
  1. (1)
    If the groups in $\mathcal{A}$ are finitely generated and accessible and there exists a JSJ tree $T$ of $G$ over $\mathcal{A}$, then there exists a JSJ tree $T^{\prime}$ of $G$ over $\mathcal{A}$ such that the halfspaces of $G\curvearrowright T^{\prime}$ are one-ended.
  2. (2)
    If $T$ is a JSJ tree of $G$ over $\mathcal{A}$ with finitely generated edge stabilizers, such that no vertex stabilizer of $T$ fixes an edge in $T$, then the halfspaces of $G\curvearrowright T$ are one-ended.
Although the definition of halfspaces for a group splitting is new, there are similar ideas that have been used many times in the literature, especially regarding the construction of JSJ trees. For example, if $(G,\mathbb{P})$ is a relatively hyperbolic group with connected Bowditch boundary, then the construction of the exact cut pair/cut point tree $T$ for $G$ [HH23] (which is the same as the JSJ tree of cylinders for elementary splittings relative to $\mathbb{P}$) can be formulated in terms of halfspaces. Indeed, the edges of $T$ correspond to decompositions of the Bowditch boundary $\partial(G,\mathbb{P})=U_{1}\cup U_{2}$ (which can be thought of as halfspaces in the boundary) such that $U_{1}\cap U_{2}$ is a cut point or an inseparable exact cut pair and $U_{1}-U_{2}$ is one of the connected components of $\partial(G,\mathbb{P})-U_{1}\cap U_{2}$. Moreover, it can be deduced from [HH23, Proposition 6.7] that if an edge $e$ in $T$ corresponds to a decomposition $\partial(G,\mathbb{P})=U_{1}\cup U_{2}$, then $U_{1}$ and $U_{2}$ are the limit sets of the halfspaces of $G\curvearrowright T$ associated to $e$ (see Definition 2). Halfspaces are also fundamental in the theory of CAT(0) cube complexes and the study of group actions on CAT(0) cube complexes – which can be regarded as a generalization of group splittings.
The next few results concern the simple connectivity at infinity and the second cohomology of groups, and their connections to one-ended halfspaces. We start with a theorem that works with an analogue of one-ended halfspaces for CW-complexes.
Theorem 1.4.
Suppose $X$ is a locally finite CW-complex, and $X_{1}$ and $X_{2}$ are connected one-ended subcomplexes of $X$ such that $X_{1}\cup X_{2}=X$. If $K$ is a finite subcomplex of $X$ (possibly empty) such that $(X_{1}\cap X_{2})-K$ has more than one unbounded component, then $X$ does not have pro-finite first homology at infinity. In particular, $X$ is not simply connected at infinity.
We note that if a finitely presented group $G$ is simply connected at infinity then $H^{2}(G,\mathbb{Z}G)=\{0\}$, and $H^{2}(G,\mathbb{Z}G)=\{0\}$ if and only if the first homology at infinity of $G$ is pro-finite (see Section §5 for the relevant definitions and results).
As a corollary of Theorem 1.4, we obtain the following generalization of Jackson’s result, working with one-ended halfspaces rather than one-ended vertex groups. Given a splitting of a one-ended group, if all the vertex groups are one-ended then all the halfspaces are also one-ended (see Lemma 2.5). The details for how we deduce Corollary 1.5 from Theorem 1.4 are given in Section 6.
Corollary 1.5.
Let $G\curvearrowright T$ be a non-trivial splitting with minimal action, with $G$ one-ended and finitely presented. Suppose the edge stabilizers are finitely generated, and suppose there is some edge stabilizer $G_{e}$ with more than one end. If the two halfspaces of $G\curvearrowright T$ associated to $e$ are one-ended, then $H^{2}(G,\mathbb{Z}G)\neq\{0\}$.
Combining Theorem 1.1 with Corollary 1.5, we get a further corollary about splittings over virtually free groups.
Corollary 1.6.
If $G$ is a finitely presented one-ended group that splits non-trivially over a virtually free group, then $H^{2}(G,\mathbb{Z}G)\neq\{0\}$.
The idea of an “artificial” splitting can be used to transform a non-trivial splitting over a subgroup $C$ into a non-trivial splitting over a larger group. We use artificial splittings to produce an example (see Example 7.2) of a finitely presented one-ended simply connected at infinity group $G$ and a non-trivial splitting of $G$ over an infinite-ended group. So, it is important to know when halfspaces are one-ended.
Perhaps the most important application of the notion of simple connectivity at infinity is to manifolds. The notion of simple connectivity at infinity was popularized by J. Stallings [Sta62], J. Munkres [Mun60], E. Luff [Luf67] and L. Siebenmann [Sie68] in the 1960’s as a way of deciding when a contractible open $n$-manifold is homeomorphic to $\mathbb{R}^{n}$ for $n\geq 5$. Following the work of Freedman in dimension 4 (see Corollary1.2 of [Fre82]) and Perelman in dimension 3 (see L. Husch and T. Price’s paper [HP70] and the addendum [HP71]), these results were extended to all dimensions $n>2$. So for contractible open topological, PL and differential manifolds $M$ of dimension $n>2$ we have that $M$ is homeomorphic to $\mathbb{R}^{n}$ if and only if $M$ is simply connected at infinity. (Note that in dimension 4 there are exotic PL and differential structures on $\mathbb{R}^{4}$ see S. Donaldson’s paper [Don83].) M. Davis [Dav83] produced closed manifolds with infinite fundamental group in dimensions $n\geq 4$ whose universal covers were contractible, and he proved these universal covers are not simply connected at infinity (and hence not homeomorphic to $\mathbb{R}^{n}$). The fundamental group $G$ of a Davis manifold is a subgroup of finite index of a right angled Coxeter group $W$ and so many “visual” splittings of $W$ (and $G$) are easily obtained. However, we have a restriction on splittings of such groups by the following corollary of Corollary 1.5.
Corollary 1.7.
If $G$ is the fundamental group of a closed aspherical $n$-manifold (or more generally an $n$-dimensional duality group, see Section 8) with $n\geq 3$, then there cannot be a splitting of $G$ over a group with more than one end where all the halfspaces are one-ended.
The contrapositive of the corollary is also interesting: if a group $G$ of cohomological dimension $n\geq 3$ splits over an infinite-ended group with all of the halfspaces one-ended, then $G$ is not a duality group. We give some examples of such groups in Section 8 (which are carefully chosen so that we cannot see another way of showing that these groups are not duality groups).
§2 contains three definitions for the halfspaces associated to a group splitting, along with a lemma to explain why they are equivalent (up to quasi-isometry).
§3 contains background about pocsets, which is needed for the proofs of Theorems 1.1 and 1.3.
§4 contains the proofs of Theorems 1.1 and 1.3, both of which are deduced from Theorem 4.1. The latter theorem essentially says that if a splitting of a one-ended group contains a halfspace with more than one-end, then one can obtain a new splitting with smaller edge stabilizers. The basic idea behind the proof is to chop up the halfspace with more than one end. From this we build a certain pocset, and the new splitting is obtained from the cubing of this pocset (which we show is a tree).
§5 begins with the definitions and connections between simple connectivity at infinity and various notions of first homology at infinity for topological spaces. Classical results show that these notions are preserved under any quasi-isometry or proper 2-equivalence between CW-complexes and can be extended to finitely presented groups.
§6 contains the proofs of Theorem 1.4 and Corollary 1.5. The proof of Theorem 1.4 is topological. We point out that it is a rare situation when $H^{2}(G,\mathbb{Z}G)$ is shown to be non-trivial by non-homological methods.
§7 begins with the notion of an artificial splitting of a group. We prove a lemma that describes how to alter a given non-trivial amalgamated product to obtain a non-trivial one with larger edge and vertex groups. Example 7.2 begins with a finitely presented group $G$ that is simply connected at infinity and splits as an amalgamated product with simply connected at infinity vertex groups and a one-ended edge group. An artificial splitting is constructed for $G$ with an infinite-ended vertex group and an infinite-ended edge group. So simply connected at infinity groups can be split non-trivially over infinite-ended groups. Corollary 1.5 implies a halfspace for this last splitting is not one-ended (so halfspaces for one-ended simply connected at infinity groups need not always be one-ended).
§8 developes two general yet elementary constructions of HNN extensions and amalgams with one-ended halfspaces (Propositions 8.1 and 8.2). We then use these constructions to produce examples of HNN extensions and amalgams that are not duality groups, but which cannot be distinguished from duality groups using previous literature (to the best of our knowledge).
2 Group splittings and their halfspaces
Definition 1. A splitting of a group $G$ is an action $G\curvearrowright T$ on a tree without edge inversions. The action is minimal if $T$ has no proper invariant subtree. The splitting is non-trivial if $G\curvearrowright T$ has no fixed point and finite if the action $G\curvearrowright T$ is cocompact. Note that any splitting of a finitely generated group with minimal action is finite.
Notation 2.1.
Whenever we talk about a graph $X$ (which will often be a tree), we denote the vertex and edge sets by $VX$ and $EX$ respectively.
We now give three definitions for the halfspaces associated to a splitting of a finitely generated group $G\curvearrowright T$ with minimal action and finitely generated edge stabilizers, along with a lemma to explain why they are equivalent. A halfspace is always a connected subgraph of a larger graph, and is considered as a metric space with respect to its induced path metric. The first two definitions of halfspace are shorter and more intuitive, while the third is more convenient for the proof of Theorem 1.1. Before giving the definitions, we note the following lemma, which implies that vertex stabilizers are always finitely generated when we have a splitting with halfspaces.
Lemma 2.2.
[Coh89, Lemma 8.32]
Any finite splitting $G\curvearrowright T$ of a finitely generated group with finitely generated edge stabilizers has finitely generated vertex stabilizers.
Definition 2.(First definition of halfspaces)
Let $G=A*_{C}B$ be a non-trivial amalgam with $A,B$ and $C$ finitely generated. Given finite generating sets $S_{A}$ and $S_{B}$ for $A$ and $B$, we know that $S=S_{A}\cup S_{B}$ is a finite generating set for $G$. Let $\operatorname{Cay}(G,S)$ denote the Cayley graph of $G$ with respect to $S$. Define $\mathfrak{h}_{A}$ (resp. $\mathfrak{h}_{B}$) to be the induced subgraph of $\operatorname{Cay}(G,S)$ whose vertices are the elements of $G$ whose normal form starts with an element of $A$ (resp. an element of $B$). We refer to $\mathfrak{h}_{A},\mathfrak{h}_{B}$ and all of their $G$-translates as halfspaces of the splitting $G=A*_{C}B$. Note that these halfspaces are connected subgraphs of $\operatorname{Cay}(G,S)$, and $\mathfrak{h}_{A}\cap\mathfrak{h}_{B}$ is the induced subgraph of $C$.
A similar definition can be made for HNN extensions, as follows. Suppose $G=A*_{C}$ is an HNN extension, with $A$ and $C$ finitely generated. Let $S_{A}$ be a finite generating set for $A$, and let $t$ be the stable letter of the HNN extension. Let $\operatorname{Cay}(G,S)$ denote the Cayley graph of $G$ with respect to $S=S+A\cup\{t\}$. Define the halfspace $\mathfrak{h}^{+}$ to be the induced subgraph of $\operatorname{Cay}(G,S)$ whose vertices are elements of $G$ with normal forms that do not start with an element of $C$ followed by a negative power of $t$. Define the halfspace $\mathfrak{h}^{-}$ to be the induced subgraph of $\operatorname{Cay}(G,S)$ whose vertices are either elements of $C$ or elements of $G$ with normal forms that do start with an element of $C$ followed by a negative power of $t$. We refer to $\mathfrak{h}^{+},\mathfrak{h}^{-}$ and all of their $G$-translates as halfspaces of the splitting $G=A*_{C}$. Note that these halfspaces are connected subgraphs of $\operatorname{Cay}(G,S)$, and $\mathfrak{h}^{+}\cap\mathfrak{h}^{-}$ is the induced subgraph of $C$.
Remark 2.3.
In Definition 2, if one chooses the finite generating set $S$ to contain a generating set for $C$, then $\mathfrak{h}_{A}\cap\mathfrak{h}_{B}$ (or $\mathfrak{h}^{+}\cap\mathfrak{h}^{-}$ in the HNN case) is connected and is isomorphic to a Cayley graph of $C$.
The second and third definitions of halfspace require the notion of halfspaces in a tree.
Definition 3. Let $T$ be a tree. Denote the midpoint of an edge $e$ by $\hat{e}$. Define a halfspace of $T$ to be the union of $\hat{e}$ with one of the components of $T-\hat{e}$. Let $\mathcal{H}(T)$ denote the set of halfspaces of $T$.
Definition 4.(Second definition of halfspaces)
Let $G\curvearrowright T$ be a splitting of a finitely generated group $G$ with minimal action and finitely generated edge stabilizers. Fix an orbit map $f:G\to VT$, and a Cayley graph $\operatorname{Cay}(G,S)$ of $G$ with respect to a finite generating set $S$. Define a halfspace of $G\curvearrowright T$ (with respect to $S$) to be an induced subgraph of $\operatorname{Cay}(G,S)$ with vertex set a preimage $f^{-1}(\mathfrak{h})$ of a halfspace $\mathfrak{h}\in\mathcal{H}(T)$. If $\mathfrak{h}$ is bounded by the edge midpoint $\hat{e}$ then we say that $f^{-1}(\mathfrak{h})$ is a halfspace of $G\curvearrowright T$ associated to $e$. For an appropriate choice of $S$, all the halfspaces of $G\curvearrowright T$ will be connected subgraphs of $\operatorname{Cay}(G,S)$. (If $S_{R}$ is the $f$-preimage of the $R$-ball in $T$ centered at $f(1)$, then one can show that all halfspaces of $G\curvearrowright T$ with respect to $S_{R}$ will be connected for large enough $R$ (extending the definition of halfspace to infinite generating sets); and if a finite generating set $S$ for $G$ contains a generating set for $G_{f(1)}$ and satisfies $S_{R}\subseteq G_{f(1)}S$, then one can show that all halfspaces of $G\curvearrowright T$ with respect to $S$ will be connected.)
Definition 5.(Third definition of halfspaces)
Let $G\curvearrowright T$ be a splitting of a finitely generated group $G$ with minimal action and finitely generated edge stabilizers. A tree of spaces for $G\curvearrowright T$ is a locally finite graph $X$ equipped with a geometric action of $G$, and a surjective $G$-equivariant map $p:X\to T$, such that:
  1. (1)
    $p$ is simplicial with respect to the first barycentric subdivision of $T$, and
  2. (2)
    the $p$-preimages of vertices and edge midpoints in $T$ are connected.
A tree of spaces can be constructed by modifying a Cayley graph of $G$, for instance apply [MSSW23, Proposition 8.2] to $G$ with the discrete topology. Define a halfspace of $X$ to be the preimage $p^{-1}(\mathfrak{h})$ of a halfspace $\mathfrak{h}\in\mathcal{H}(T)$. Write $\mathcal{H}(X)$ for the set of halfspaces of $X$. Note that halfspaces of $X$ are connected subgraphs of $X$.
We now prove that these three definitions of halfspace are well-defined and equivalent up to quasi-isometry. This is not as straightforward to prove as one might imagine, since the inclusion of a halfspace $\mathfrak{h}\xhookrightarrow{}\operatorname{Cay}(G,S)$ (say for the second definition) is not necessarily a quasi-isometric embedding.
Lemma 2.4.
The three definitions for halfspaces given above are coarsely well-defined and coarsely equivalent. More precisely, if we fix a splitting $G\curvearrowright T$ and a halfspace $\mathfrak{h}$ of $T$, then the halfspace of $G\curvearrowright T$ corresponding to $\mathfrak{h}$ remains the same up to quasi-isometry when we change the choice of finite generating set for $G$ and when we switch between the three definitions of halfspace (and when we change the orbit map $f$ from the second definition and the graph $X$ from the third definition). In particular, the notion of the halfspaces of a splitting $G\curvearrowright T$ being one-ended is well-defined.
Proof.
We begin by showing that the first definition is a special case of the second. Let $G=A*_{C}B$ be a non-trivial amalgam with $A,B$ and $C$ finitely generated, and let $G\curvearrowright T$ be the corresponding Bass–Serre tree. Let $v_{A},v_{B}\in VT$ be the vertices stabilized by $A$ and $B$ respectively, and let $e$ be the edge joining $v_{A}$ and $v_{B}$. Let $\mathfrak{h}\subseteq T$ be the halfspace bounded by $\hat{e}$ that contains $v_{A}$. Let $f:G\to VT$ be the orbit map given by $g\mapsto gv_{A}$ and let $S=S_{A}\cup S_{B}$ be the finite generating set for $G$ from the first definition. Then it is not hard to show that the halfspace $\mathfrak{h}_{A}$ from the first definition is exactly the same as the halfspace from the second definition with vertex set $f^{-1}(\mathfrak{h})$ (with respect to generating set $S$). Similar arguments for $\mathfrak{h}_{B}$ and for the $G$-translates of $\mathfrak{h}_{A}$ and $\mathfrak{h}_{B}$ show that the set of halfspaces from the first definition is the same as the set of halfspaces from the second definition. A similar argument works for HNN extensions.
Next, we show the equivalence of the second and third definitions. Let $e\in ET$ and let $\mathfrak{h}\in\mathcal{H}(T)$ be one of the halfspaces bounded by $\hat{e}$. We have an orbit map $f:G\to T$ from the second definition and a tree of spaces $p:X\to T$ from the third definition. Let $i:G\to VX$ be an orbit map for the action of $G$ on $X$ such that the following diagram commutes.
${G}$${X}$${T}$$\scriptstyle{f}$$\scriptstyle{i}$$\scriptstyle{p}$ (1)
The second definition provides a halfspace $\mathfrak{h}_{1}\subseteq\operatorname{Cay}(G,S)$ (with respect to a suitable finite generating set $S$) with vertex set $V\mathfrak{h}_{1}=f^{-1}(\mathfrak{h})$, while the third definition provides a halfspace $\mathfrak{h}_{2}:=p^{-1}(\mathfrak{h})\subseteq X$. Denote the metrics on $\operatorname{Cay}(G,S)$ and $X$ by $d_{S}$ and $d_{X}$ respectively, and denote the induced path metrics on $\mathfrak{h}_{1}$ and $\mathfrak{h}_{2}$ by $d_{\mathfrak{h}_{1}}$ and $d_{\mathfrak{h}_{2}}$ respectively. Observe that $i(V\mathfrak{h}_{1})\subseteq V\mathfrak{h}_{2}$. We claim that the map $i:(V\mathfrak{h}_{1},d_{\mathfrak{h}_{1}})\to(V\mathfrak{h}_{2},d_{\mathfrak{h}_{2}})$ is a quasi-isometry.
The Milnor–Svarc Lemma implies that $i:(G,d_{S})\to(X,d_{X})$ is a quasi-isometry. Say $C\geq 1$ is such that
$C^{-1}d_{S}(g,g^{\prime})-C\leq d_{X}(i(g),i(g^{\prime}))\leq Cd_{S}(g,g^{\prime})+C$ (2)
for all $g,g^{\prime}\in G$, and such that the $C$-neighborhood of $i(G)$ is equal to the whole of $X$.
Let $\hat{\mathfrak{h}}_{2}:=p^{-1}(\hat{e})\subseteq\mathfrak{h}_{2}$. Since $G_{e}$ acts cocompactly on the $4C$-neighborhood of $\hat{\mathfrak{h}}_{2}$, preserving $\mathfrak{h}_{2}$, there is a constant $L$ such that, for any $v,v^{\prime}\in V\mathfrak{h}_{2}$, if $d_{X}(v,v^{\prime})\leq 2C$ and $d_{X}(v,\hat{\mathfrak{h}}_{2})\leq 2C$ then $d_{\mathfrak{h}_{2}}(v,v^{\prime})\leq L$. We may also assume $L\geq 2C$. Now consider adjacent vertices $g,g^{\prime}\in V\mathfrak{h}_{1}$. By (2), we have $d_{X}(i(g),i(g^{\prime}))\leq Cd_{S}(g,g^{\prime})+C=2C$. Note that any path leaving $\mathfrak{h}_{2}$ must cross $\hat{\mathfrak{h}}_{2}$, so if $d_{X}(i(g),\hat{\mathfrak{h}}_{2})\geq 2C$ then $i(g^{\prime})\in\mathfrak{h}_{2}$ and $d_{\mathfrak{h}_{2}}(i(g),i(g^{\prime}))=d_{X}(i(g),i(g^{\prime}))\leq 2C$. Otherwise, if $d_{X}(i(g),\hat{\mathfrak{h}}_{2})<2C$ then $d_{\mathfrak{h}_{2}}(i(g),i(g^{\prime}))\leq L$ by definition of $L$. In either case, we have $d_{\mathfrak{h}_{2}}(i(g),i(g^{\prime}))\leq L$. By mapping paths in $\mathfrak{h}_{1}$ into $X$, we deduce that
$d_{\mathfrak{h}_{2}}(i(g),i(g^{\prime}))\leq Ld_{\mathfrak{h}_{1}}(g,g^{\prime})$ (3)
for all $g,g^{\prime}\in V\mathfrak{h}_{1}$.
Since $G_{e}$ acts cocompactly on the $C$-neighborhood of $\hat{\mathfrak{h}}_{2}$, and $G_{e}$ stabilizes $\mathfrak{h}_{1}$, there is a constant $K$ such that, for any $v\in V\mathfrak{h}_{2}$, if $d_{X}(v,\hat{\mathfrak{h}}_{2})\leq C$ then $d_{\mathfrak{h}_{2}}(v,i(V\mathfrak{h}_{1}))\leq K$. We may also assume $K\geq C$. We now claim that
$d_{\mathfrak{h}_{2}}(v,i(V\mathfrak{h}_{1}))\leq K$ (4)
for all $v\in V\mathfrak{h}_{2}$. Indeed, if $d_{X}(v,\hat{\mathfrak{h}}_{2})\leq C$ then the claim follows by definition of $K$. Otherwise, if $d_{X}(v,\hat{\mathfrak{h}}_{2})>C$, then take $g\in G$ with $d_{X}(v,i(g))\leq C$ (which exists by definition of $C$). The shortest path from $v$ to $i(g)$ does not cross $\hat{\mathfrak{h}}_{2}$, so it stays within $\mathfrak{h}_{2}$, which means $g\in\mathfrak{h}_{1}$ and $d_{\mathfrak{h}_{1}}(v,i(g))=d_{X}(v,i(g))\leq C\leq K$.
Put $D=2C^{2}(3K+1+C)+K$. Using properness of the $G$-action on $X$, and the fact that $G_{e}$ acts cocompactly on the $(D+K)$-neighborhood of $\hat{\mathfrak{h}}_{2}$, there is a constant $M$ such that, for any $g,g^{\prime}\in V\mathfrak{h}_{1}$, if $d_{X}(i(g),\hat{\mathfrak{h}}_{2})\leq D+K$ and $d_{X}(i(g),i(g^{\prime}))\leq 2K+1$ then $d_{\mathfrak{h}_{1}}(g,g^{\prime})\leq M$. And we may assume $M\geq D$. Now take $g,g^{\prime}\in V\mathfrak{h}_{1}$. Let $i(g)=v_{0},v_{1},\dots,v_{n}=i(g^{\prime})$ be the vertices of a shortest path in $\mathfrak{h}_{2}$ from $i(g)$ to $i(g^{\prime})$. For each $v_{j}$, take $g_{j}\in V\mathfrak{h}_{1}$ with $d_{\mathfrak{h}_{1}}(v_{j},i(g_{j}))\leq K$ (which exists by (4)). We may assume that $g_{0}=g$ and $g_{n}=g^{\prime}$. Now consider a pair of consecutive vertices $v_{j},v_{j+1}$. Note that
$d_{X}(i(g_{j}),i(g_{j+1}))\leq d_{X}(i(g_{j}),v_{j})+d_{X}(v_{j},v_{j+1})+d_{X}(v_{j+1},i(g_{j+1}))\leq 2K+1.$ (5)
If one of $v_{j},v_{j+1}$ is in the $D$-neighborhood of $\hat{\mathfrak{h}}_{2}$, then one of $i(g_{j}),i(g_{j+1})$ is in the $(D+K)$-neighborhood of $\hat{\mathfrak{h}}_{2}$, so $d_{\mathfrak{h}_{1}}(g_{j},g_{j+1})\leq M$ by (5) and definition of $M$. Now suppose neither $v_{j}$ nor $v_{j+1}$ is in the $D$-neighborhood of $\hat{\mathfrak{h}}_{2}$. By (2) and (5) we have $d_{S}(g_{j},g_{j+1})\leq C(2K+1+C)$. Moreover, any shortest path $\gamma$ in $\operatorname{Cay}(G,S)$ from $g_{j}$ to $g_{j+1}$ will have $i$-image contained in the $2C^{2}(3K+1+C)$-neighborhood of $i(g_{j})$ (again using (2)), which is contained in $\mathfrak{h}_{2}$ since
$d_{X}(i(g_{j}),\hat{\mathfrak{h}}_{2})\geq d_{X}(v_{j},\hat{\mathfrak{h}}_{2})-d_{X}(v_{j},i(g_{j}))>D-K=2C^{2}(3K+1+C).$
So $\gamma\subseteq\mathfrak{h}_{1}$ and
$d_{\mathfrak{h}_{1}}(g_{j},g_{j+1})=d_{S}(g_{j},g_{j+1})\leq C(2K+1+C)<D\leq M.$
In either case we get $d_{\mathfrak{h}_{1}}(g_{j},g_{j+1})\leq M$, hence
$d_{\mathfrak{h}_{1}}(g,g^{\prime})\leq Mn=Md_{\mathfrak{h}_{2}}(i(g),i(g^{\prime})).$ (6)
The combination of (3), (4) and (6) tells us that $i:(V\mathfrak{h}_{1},d_{\mathfrak{h}_{1}})\to(V\mathfrak{h}_{2},d_{\mathfrak{h}_{2}})$ is a quasi-isometry, as required. Thus the halfspaces from the second definition will be quasi-isometric to the halfspaces from the third definition. We can also deduce from this that the halfspaces from both the second and third definitions are coarsely well-defined – for instance, if $S^{\prime}$ is another finite generating set for $G$, and $\mathfrak{h}^{\prime}_{1}\subseteq\operatorname{Cay}(G,S^{\prime})$ is the halfspace with vertex set $V\mathfrak{h}^{\prime}_{1}=f^{-1}(\mathfrak{h})$, then the above argument tells us that the halfspaces $\mathfrak{h}_{1}$ and $\mathfrak{h}^{\prime}_{1}$ are both quasi-isometric to $\mathfrak{h}_{2}$, hence they are quasi-isometric to each other. ∎
We finish this section with a lemma about the connection between one-endedness of vertex stabilizers and one-endedness of halfspaces.
Lemma 2.5.
Let $G\curvearrowright T$ be a splitting of a one-ended group $G$ with minimal action and finitely generated edge stabilizers. Let $\mathfrak{h}\subseteq T$ be a halfspace bounded by an edge midpoint $\hat{e}$, and let $v\in VT$ be the unique vertex in $\mathfrak{h}$ incident to $e$. If $G_{v}$ is one-ended then the halfspace of $G\curvearrowright T$ corresponding to $\mathfrak{h}$ is one-ended.
Proof.
As in Definition 2, let $p:X\to T$ be a tree of spaces for $G\curvearrowright T$. We must show that $p^{-1}(\mathfrak{h})$ is one-ended, so suppose for contradiction that $K\subseteq p^{-1}(\mathfrak{h})$ is a finite subgraph such that $p^{-1}(\mathfrak{h})$ has multiple unbounded components. We are assuming that $G$ and $G_{v}$ are one-ended, hence $X$ and $p^{-1}(v)$ are one-ended as well. In particular $p^{-1}(v)-K$ only has one unbounded component, call it $C_{1}$. Now consider the finite subgraph
$K^{\prime}:=K\cup(p^{-1}(v)-C_{1})\subseteq X.$
Let $C_{2}$ be an unbounded component of $p^{-1}(\mathfrak{h})-K$ disjoint from $C_{1}$. Observe that $p^{-1}(\hat{e})$ is contained in a bounded neighborhood of $p^{-1}(v)$ with respect to the induced metric of $p^{-1}(\mathfrak{h})$, hence also in a bounded neighborhood of $C_{1}$. It follows that $C_{2}\cap p^{-1}(\hat{e})$ is bounded, so only finitely many vertices in $C_{2}$ lie in $p^{-1}(\hat{e}\cup v)$. Any vertex in $C_{2}-p^{-1}(\hat{e}\cup v)$ is separated from $C_{1}$ in $X$ by $K^{\prime}$, so this contradicts one-endedness of $X$. ∎
3 Pocsets
We now recall the construction of a CAT(0) cube complex from a pocset or wallspace. This construction was originally due to Sageev [Sag95], although our formulation will be closer to the version in [Man20] – see also [Nic04, CN05, Rol16, Wis12].
Definition 6. A pocset is a poset $(\mathcal{P},\leq)$ together with an involution $A\mapsto A^{*}$ for all $A\in\mathcal{P}$ satisfying:
  1. (1)
    $A$ and $A^{*}$ are incomparable.
  2. (2)
    $A\leq B\Rightarrow B^{*}\leq A^{*}$.
Elements $A,B\in\mathcal{P}$ are nested if $A\leq B$ or $B\leq A$. And $A,B\in\mathcal{P}$ are transverse if neither $A,B$ nor $A,B^{*}$ are nested. The width of $\mathcal{P}$ is the maximum number of pairwise transverse elements, if such a maximum exists, otherwise we say the width is $\infty$.
Definition 7. An ultrafilter on a pocset $\mathcal{P}$ is a subset $\omega\subseteq\mathcal{P}$ satisfying:
  1. (1)
    (Completeness) For every $A\in\mathcal{P}$, exactly one of $\{A,A^{*}\}$ is in $\omega$.
  2. (2)
    (Consistency) If $A\in\omega$ and $A\leq B$, then $B\in\omega$.
An ultrafilter $\omega$ is DCC (descending chain condition) if it contains no strictly descending infinite chain $A_{1}>A_{2}>A_{3}>\cdots$.
Proposition 3.1.
Let $\mathcal{P}$ be a pocset of finite width that admits at least one DCC ultrafilter. Then there is a CAT(0) cube complex $C=C(\mathcal{P})$, called the cubing of $\mathcal{P}$, such that:
  1. (1)
    The vertices of $C$ are the DCC ultrafilters on $\mathcal{P}$.
  2. (2)
    Two vertices $\omega_{1},\omega_{2}$ in $C$ are joined by an edge if and only if $\omega_{1}\triangle\omega_{2}=\{A,A^{*}\}$ for some $A\in\mathcal{P}$.
  3. (3)
    The dimension of $C$ is equal to the width of $\mathcal{P}$.
In this paper we will only deal with pocsets whose cubings are trees, or equivalently, pocsets that contain no pairs of transverse elements. One source of pocsets will come from the following definition.
Definition 8. A wallspace $(X,\mathcal{P})$ is a set $X$ together with a family $\mathcal{P}$ of non-empty subsets that is closed under complementation, such that for any $x,y\in X$ the set $\{A\in\mathcal{P}\mid x\in A,\,y\notin A\}$ is finite. $\mathcal{P}$ forms a pocset under inclusion, with the involution given by complementation. Moreover, for any $x\in X$ the set
$\lambda(x):=\{A\in\mathcal{P}\mid x\in A\}$
is a DCC ultrafilter. Therefore, if $\mathcal{P}$ has finite width, we can form the cubing $C=C(\mathcal{P})$, and we have a map $\lambda:X\to C$.
Another pocset that will frequently be used in this paper is the pocset of halfspaces of a tree, or a tree of spaces. These pocsets are not wallspaces according to Definition 3, although they are very similar.
Definition 9. Let $T$ be a tree. Recall the definition of halfspaces in $T$ (Definition 2), and let $\mathcal{H}(T)$ denote the set of halfspaces. Let $\mathfrak{h}\mapsto\mathfrak{h}^{*}$ be the involution of $\mathcal{H}(T)$ that exchanges the complementary components of each edge midpoint $\hat{e}$. This makes $\mathcal{H}(T)$ into a pocset, with partial order given by inclusion, and the cubing of this pocset recovers the tree $T$.
Definition 10. Let $p:X\to T$ be a tree of spaces associated to a group splitting $G\curvearrowright T$ (Definition 2). Let $\mathcal{H}(X)$ denote the set of halfspaces of $X$ (recall this is just the set of $p$-preimages of halfspaces of $T$), and define an involution $\mathfrak{h}\mapsto\mathfrak{h}^{*}$ on $\mathcal{H}(X)$ so that $p(\mathfrak{h}^{*})=p(\mathfrak{h})^{*}$. The partial order of inclusion makes $\mathcal{H}(X)$ into a pocset, and $p$ induces a pocset isomorphism between $\mathcal{H}(X)$ and $\mathcal{H}(T)$. For $\mathfrak{h}\in\mathcal{H}(X)$, let $\hat{\mathfrak{h}}:=\mathfrak{h}\cap\mathfrak{h}^{*}$, and call this a wall. Note that walls are precisely the $p$-preimages of edge midpoints in $T$. Note that halfspaces and walls in $X$ are connected subgraphs of $X$, and the walls are pairwise disjoint.
Finally, we give the following lemma, to be used in Section 4.
Lemma 3.2.
Let $p:X\to T$ be a tree of spaces associated to a group splitting $G\curvearrowright T$. For each $\mathfrak{h}\in\mathcal{H}(X)$, we have that $\mathfrak{h}$ is not contained in a bounded neighborhood of $\hat{\mathfrak{h}}$ or $\mathfrak{h}^{*}$. Moreover, if $G$ (or equivalently $X$) is one-ended, then $\hat{\mathfrak{h}}$ is unbounded.
Proof.
The action $G\curvearrowright T$ is minimal, so $p(\mathfrak{h})$ is not contained in a bounded neighborhood of $p(\hat{\mathfrak{h}})$ or $p(\mathfrak{h}^{*})$. Since $p$ is distance non-increasing, we have that $\mathfrak{h}$ is not contained in a bounded neighborhood of $\hat{\mathfrak{h}}$ or $\mathfrak{h}^{*}$. Secondly, as $\hat{\mathfrak{h}}$ separates $\mathfrak{h}-\hat{\mathfrak{h}}$ from $\mathfrak{h}^{*}-\hat{\mathfrak{h}}$ in $X$, and since $X$ is one-ended, we must have that $\hat{\mathfrak{h}}$ is unbounded. ∎
4 Chopping up halfspaces
In this section we prove Theorems 1.1 and 1.3. The main technical result we need is the following.
Theorem 4.1.
Let $G\curvearrowright T$ be a splitting with minimal action and finitely generated edge stabilizers, and with $G$ one-ended and finitely generated. If the splitting has some halfspace with more than one end, then there is another splitting $G\curvearrowright T^{\prime}$ with minimal action and finitely generated edge stabilizers, such that every edge stabilizer (resp. vertex stabilizer) of $T^{\prime}$ is contained in an edge stabilizer (resp. vertex stabilizer) of $T$, and for each $e\in ET$ there is a tree $T_{e}$ and there is an action of $G$ on $\sqcup_{e\in ET}T_{e}$ that is compatible with the action on $T$ such that:
  1. (a)
    For each $e\in ET$, either $T_{e}$ is a single vertex or the action $G_{e}\curvearrowright T_{e}$ is a non-trivial finite splitting with finite edge stabilizers.
  2. (b)
    There is an edge $e_{0}\in ET$ such that $T_{e_{0}}$ is not a single vertex.
  3. (c)
    There is an injective $G$-equivariant map $ET^{\prime}\to\sqcup_{e}VT_{e}$.
  4. (d)
    If no vertex stabilizer of $T$ fixes an edge in $T$, then there is a vertex $v_{0}\in VT$ such that $G_{v_{0}}$ has no fixed point in $T^{\prime}$.
Before proving Theorem 4.1, let’s see how to deduce Theorems 1.1 and 1.3. This requires the following theorem concerning accessible groups.
Theorem 4.2.
[Dic80, Theorem 5.12]
Let $G$ be a finitely generated accessible group and let $(G_{i})$ be a sequence of groups such that $G_{0}=G$, and $G_{i+1}$ is a vertex group in some non-trivial finite splitting of $G_{i}$ over finite subgroups. Then the sequence $(G_{i})$ terminates.
Theorem 1.1.
Let $G\curvearrowright T$ be a non-trivial splitting with $G$ one-ended and finitely generated. Suppose the edge stabilizers are finitely generated and accessible. Then there is a non-trivial splitting $G\curvearrowright T^{\prime}$ with minimal action such that:
  1. (1)
    The halfspaces of $G\curvearrowright T^{\prime}$ are one-ended.
  2. (2)
    Edge stabilizers (resp. vertex stabilizers) of $T^{\prime}$ are finitely generated and are subgroups of the edge stabilizers (resp. vertex stabilizers) of $T$.
  3. (3)
    For each edge $e^{\prime}$ in $T^{\prime}$ there exists an edge $e$ in $T$ such that $G_{e^{\prime}}$ is a vertex stabilizer in some finite splitting of $G_{e}$ over finite subgroups.
Proof.
Passing to a minimal subtree if necessary, we may assume that the action $G\curvearrowright T$ is minimal. If the halfspaces of the splitting are already one-ended then we are done. Otherwise, apply Theorem 4.1 repeatedly until we get a splitting where all the halfspaces are one-ended. Each time we apply Theorem 4.1, the new edge stabilizers (resp. vertex stabilizers) will be finitely generated (using Lemma 2.2), and they will be subgroups of the old edge stabilizers (resp. vertex stabilizers). Moreover, by (a) each old edge stabilizer admits a finite splitting over finite subgroups which is either an action on a single vertex or a non-trivial splitting, and at least one of these splittings is non-trivial by (b). Furthermore, the new edge stabilizers are vertex stabilizers in the splittings of the old edge stabilizers by (c). Since the original edge stabilizers are accessible, Theorem 4.2 ensures that the process of repeatedly applying Theorem 4.1 cannot go on forever, so we must eventually obtain a splitting of $G$ where all the halfspaces are one-ended. The second and third parts of Theorem 1.1 also follow readily from the construction. ∎
Definition 11. [GL17] Let $G$ be a group and let $\mathcal{A}$ be a family of subgroups of $G$ that is closed under conjugating and taking subgroups. If $G\curvearrowright T$ is a splitting with minimal action and edge stabilizers in $\mathcal{A}$, then we refer to $T$ as an $\mathcal{A}$-tree of $G$. An $\mathcal{A}$-tree $T^{\prime}$ dominates an $\mathcal{A}$-tree $T$ if every vertex stabilizer of $T^{\prime}$ fixes a point in $T$. A JSJ tree of $G$ over $\mathcal{A}$ is an $\mathcal{A}$-tree $T$ such that:
  • $T$ is universally elliptic, meaning its edge stabilizers fix points in all other $\mathcal{A}$-trees, and
  • $T$ dominates all other universally elliptic $\mathcal{A}$-trees.
Theorem 1.3.
Let $G$ be a finitely generated one-ended group and let $\mathcal{A}$ be a family of subgroups of $G$ that is closed under conjugating and taking subgroups.
  1. (1)
    If the groups in $\mathcal{A}$ are finitely generated and accessible and there exists a JSJ tree $T$ of $G$ over $\mathcal{A}$, then there exists a JSJ tree $T^{\prime}$ of $G$ over $\mathcal{A}$ such that the halfspaces of $G\curvearrowright T^{\prime}$ are one-ended.
  2. (2)
    If $T$ is a JSJ tree of $G$ over $\mathcal{A}$ with finitely generated edge stabilizers, such that no vertex stabilizer of $T$ fixes an edge in $T$, then the halfspaces of $G\curvearrowright T$ are one-ended.
Proof.
We start with a general observation about JSJ trees. Suppose $T$ is a JSJ tree of $G$ over $\mathcal{A}$, suppose $T^{\prime}$ is another $\mathcal{A}$-tree of $G$ that dominates $T$, and suppose that the edge stabilizers of $T^{\prime}$ are subgroups of the edge stabilizers of $T$. It then follows from Definition 4 that
  1. (i)
    $T^{\prime}$ is also a JSJ tree for $G$ over $\mathcal{A}$, and
  2. (ii)
    $T$ dominates $T^{\prime}$.
Let’s now show (1), so suppose the groups in $\mathcal{A}$ are finitely generated and accessible and let $T$ be a JSJ tree of $G$ over $\mathcal{A}$. By (i), we may apply Theorem 1.1 to produce a JSJ tree $T^{\prime}$ of $G$ over $\mathcal{A}$ such that the halfspaces of $G\curvearrowright T^{\prime}$ are one-ended.
Let’s now prove (2), so suppose $T$ is a JSJ tree of $G$ over $\mathcal{A}$ with finitely generated edge stabilizers, such that no vertex stabilizer of $T$ fixes an edge in $T$, and suppose for contradiction that $G\curvearrowright T$ does not have one-ended halfspaces. By (i), we may apply Theorem 4.1 to produce a JSJ tree $T^{\prime}$ of $G$ over $\mathcal{A}$. But part (d) in Theorem 4.1 implies that $T$ does not dominate $T^{\prime}$, contradicting (ii). ∎
4.1 Minimal cuts
In this subsection we recall work of Krön [Krö10] about cutting up graphs with more than one end in an automorphism-invariant way. Krön’s paper is in turn based on work of Dunwoody [Dun82].
Definition 12. Let $Y$ be a connected simplicial graph. For $C\subseteq VY$, define the edge boundary of $C$, denoted $\delta C$, to be the set of edges in $Y$ with one vertex in $C$ and one vertex in $VY-C$. A ray in $Y$ is an embedded one-way infinite path in $Y$. A cut is a subset $C\subseteq VY$ with $\delta C$ finite and such that the induced subgraphs of $C$ and $VY-C$ are both connected and contain rays. Note that, if $Y$ is locally finite, then a connected subgraph of $Y$ contains a ray if and only if it is unbounded. A cut $C$ is minimal if $|\delta C|$ is minimal.
Remark 4.3.
Any locally finite simplicial graph with more than one end has a cut.
Proposition 4.4.
Let $Y$ be a locally finite simplicial graph. Given a cut $C$, we can form a wallspace $(VY,\mathcal{P}(Y,C))$ (as in Definition 3), where
$\mathcal{P}(Y,C):=\{gC,VY-gC\mid g\in\operatorname{Aut}(Y)\}.$
Proof.
Clearly, $\mathcal{P}(Y,C)$ is a family of non-empty subsets of $VY$ that is closed under complementation. It remains to show that, for $x,y\in VY$, the set $S=\{A\in\mathcal{P}(Y,C)\mid x\in A,\,y\notin A\}$ is finite. Indeed, fix a path $\gamma$ from $x$ to $y$. Given $A\in S$, we either have $A=gC$ or $A=VY-gC$ for some $g\in\operatorname{Aut}(Y)$. Now, $\delta(gC)$ determines the pair $(gC,VY-gC)$, so it suffices to show that there are only finitely many possibilities for $\delta(gC)$. But the path $\gamma$ contains an edge in $\delta(gC)$, and $\delta(gC)$ is an $\operatorname{Aut}(Y)$-translate of the finite edge set $\delta C$, so the local finiteness of $Y$ implies that there are only finitely many possibilities for $\delta(gC)$. ∎
The key result of [Krö10] we will use is the following.
Theorem 4.5.
[Krö10, Theorem 3.3]
Let $Y$ be a locally finite simplicial graph with more than one end. Then $Y$ has a minimal cut $C$ such that the pocset $\mathcal{P}(Y,C)$ defined in Proposition 4.4 has no pairs of transverse elements, or, equivalently, such that $C(\mathcal{P}(Y,C))$ is a tree.
Remark 4.6.
Krön’s theorem is actually in the context of simple graphs, but one can deduce the same result for graphs that contain multi-edges. Indeed, if a graph $Y$ contains multi-edges then one can consider its first barycentric subdivision $\dot{Y}$, which is a simple graph. There is a correspondence between cuts in $Y$ and cuts in $\dot{Y}$, so if the theorem holds for $\dot{Y}$ it must also hold for $Y$.
4.2 Chopping up a halfspace $\mathfrak{h}_{0}$
We now turn to the proof of Theorem 4.1, which will be spread over Subsections 4.24.5. Let $G\curvearrowright T$ be a splitting as in the theorem, and suppose that it does not have one-ended halfspaces. Let $p:X\to T$ be a tree of spaces associated to $G\curvearrowright T$ (Definition 2).
Define a halfspace cut to be a pair $(\mathfrak{h},C)$ where $\mathfrak{h}\in\mathcal{H}(X)$ and $C$ is a cut in $\mathfrak{h}$. As in Definition 4.1, let $\delta C$ denote the set of edges in $\mathfrak{h}$ with one vertex in $C$ and one vertex in $V\mathfrak{h}-C$. As the splitting $G\curvearrowright T$ does not have one-ended halfspaces, there exists a halfspace $\mathfrak{h}\in\mathcal{H}(X)$ that is not one-ended, and there exists $C\subseteq V\mathfrak{h}$ such that $(\mathfrak{h},C)$ is a halfspace cut.
Lemma 4.7.
If $(\mathfrak{h},C)$ is a halfspace cut, then $C\cap V\hat{\mathfrak{h}}$ and $V\hat{\mathfrak{h}}-C$ are both unbounded, and $\delta C$ contains at least one edge in $\hat{\mathfrak{h}}$.
Proof.
Suppose $C\cap V\hat{\mathfrak{h}}$ is bounded (argument similar for $V\hat{\mathfrak{h}}-C$). The wall $\hat{\mathfrak{h}}$ is unbounded by Lemma 3.2, so $V\hat{\mathfrak{h}}-C$ is unbounded. Let $\delta^{+}$ be the union of $\delta C$ and all edges in $X$ that are incident to vertices in $C\cap V\hat{\mathfrak{h}}$. Any path in $X$ that starts in $C$ and leaves $\mathfrak{h}$ must cross over $\hat{\mathfrak{h}}$, and any path in $\mathfrak{h}$ that starts in $C$ and reaches $V\mathfrak{h}-C$ must cross over $\delta C$. Hence $\delta^{+}$ is a finite set of edges that separates $C$ from $V\mathfrak{h}-C$ in $X$, contradicting one-endedness of $X$ (note that $C$ and $V\mathfrak{h}-C$ are both unbounded).
Since $C\cap\hat{\mathfrak{h}}$ and $\hat{\mathfrak{h}}-C$ are both unbounded, so in particular non-empty, there exists a path in $\hat{\mathfrak{h}}$ from $C\cap\hat{\mathfrak{h}}$ to $\hat{\mathfrak{h}}-C$ (recall that $\hat{\mathfrak{h}}$ is connected). As $\delta C$ separates $C$ from $V\mathfrak{h}-C$ in $\mathfrak{h}$, this path must contain an edge in $\delta C$. ∎
Given a halfspace cut $(\mathfrak{h},C)$, let $W(C)$ be the number of edges in $\delta C$ that are contained in walls of $X$. Say that $(\mathfrak{h},C)$ is $W$-minimal if $W(C)$ is minimal among all halfspace cuts.
Lemma 4.8.
Let $(\mathfrak{h},C)$ be a $W$-minimal halfspace cut. If $\delta C$ separates a halfspace $\mathfrak{h}^{\prime}\subsetneq\mathfrak{h}$ into more than one component, then only one of these components is unbounded.
Proof.
Suppose for contradiction that $\delta C$ separates a halfspace $\mathfrak{h}^{\prime}\subsetneq\mathfrak{h}$ into more than one unbounded component. Let $\delta^{\prime}\subseteq\delta C\cap E\mathfrak{h}^{\prime}$ be such that $\mathfrak{h}^{\prime}-\delta^{\prime}$ has exactly two unbounded components, $C^{\prime}$ and $D^{\prime}$. This means that $(\mathfrak{h}^{\prime},C^{\prime})$ is a halfspace cut, with $\delta C^{\prime}=\delta^{\prime}$. Lemma 4.7 tells us that $\delta C$ contains an edge in $\hat{\mathfrak{h}}$, and as this edge is not in $\delta C^{\prime}$ we deduce that $W(C^{\prime})<W(C)$, contradicting $W$-minimality of $(\mathfrak{h},C)$. ∎
Lemma 4.9.
Let $(\mathfrak{h},C)$ be a $W$-minimal halfspace cut. Then each halfspace $\mathfrak{h}^{\prime}\subsetneq\mathfrak{h}$ satisfies either $V\mathfrak{h}^{\prime}\subseteq C$ or $V\mathfrak{h}^{\prime}\subseteq V\mathfrak{h}-C$.
Proof.
Suppose for contradiction that $\mathfrak{h}^{\prime}\subsetneq\mathfrak{h}$ is a halfspace such that $C\cap V\mathfrak{h}^{\prime}$ and $V\mathfrak{h}^{\prime}-C$ are both non-empty. By Lemma 4.8, only one of $C\cap V\mathfrak{h}^{\prime}$ and $V\mathfrak{h}^{\prime}-C$ is infinite. Say $C\cap V\mathfrak{h}^{\prime}$ is finite. Now $C\cap V\mathfrak{h}^{\prime}$ separates the induced subgraph of $C$ into only finitely many components, so at least one of these components is unbounded – say it has vertex set $D$. As both $\mathfrak{h}^{\prime}$ and the induced subgraph of $V\mathfrak{h}-C$ are connected, we see that the induced subgraph of $V\mathfrak{h}-D$ is connected. Furthermore, the edge boundary of $D$ in $\mathfrak{h}$ is finite since it is contained in the union of $\delta C$ and the set of edges that meet $C\cap V\mathfrak{h}^{\prime}$. Hence $(\mathfrak{h},D)$ is a halfspace cut.
We now obtain a contradiction by showing that $W(D)<W(C)$. Indeed, $\delta D-\delta C$ only contains edges that join a vertex in $D\subseteq\mathfrak{h}-\mathfrak{h}^{\prime}$ to a vertex in $C\cap V\hat{\mathfrak{h}}^{\prime}$, and these edges are not contained in walls. On the other hand, any path in the induced subgraph of $C$ from $\mathfrak{h}^{\prime}$ to $C-\mathfrak{h}^{\prime}$ crosses $\hat{\mathfrak{h}}^{\prime}$, so $C\cap V\hat{\mathfrak{h}}^{\prime}\neq\emptyset$, and any path in $\hat{\mathfrak{h}}^{\prime}$ from $C\cap V\hat{\mathfrak{h}}^{\prime}$ to $V\hat{\mathfrak{h}}^{\prime}-C$ crosses $\delta C$, so $\delta C-\delta D$ contains at least one edge in the wall $\hat{\mathfrak{h}}^{\prime}$. ∎
[…]
Figure 1: Cartoon picture for the proof of Lemma 4.9.
Lemma 4.10.
Modifying $X$ if necessary, we may assume that all halfspace cuts $(\mathfrak{h},C)$ with $|\delta C|$ minimal are also $W$-minimal.
Proof.
The modification of $X$ we consider is where we fix an integer $n\geq 1$, and replace each edge in each wall of $X$ with $n$ edges that join the same pair of vertices. Let $X_{n}$ be the resulting graph, which is endowed with a collection of halfspaces and walls that are inherited from $X$ in the obvious way. Any halfspace cut $(\mathfrak{h},C)$ in $X$ is still a halfspace cut in $X_{n}$, and vice versa. The only difference is that the edge boundary $\delta C$ might become larger in $X_{n}$ because of the extra edges. More precisely, the numbers $W(C)$ and $|\delta C|$ both increase by $(n-1)W(C)$ when passing from $X$ to $X_{n}$. Note that a halfspace cut is $W$-minimal with respect to $X$ if and only if it is $W$-minimal with respect to $X_{n}$, but the property of having minimal $|\delta C|$ may change when passing between $X$ and $X_{n}$. Let $w$ be the smallest value of $W(C)$ among halfspace cuts $(\mathfrak{h},C)$ in $X$, and let $N$ be the smallest value of $|\delta C|$ among $W$-minimal halfspace cuts $(\mathfrak{h},C)$ in $X$. Put $n=N+1$. Then the smallest possible value of $|\delta C|$ in $X_{n}$ is $N+(n-1)w$, and this is only achieved by $W$-minimal halfspace cuts, as required. ∎
Now let $(\mathfrak{h}_{0},C)$ be a halfspace cut with $|\delta C|$ minimal. Lemmas 4.9 and 4.10 imply that each halfspace $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$ satisfies either $V\mathfrak{h}\subseteq C$ or $V\mathfrak{h}\subseteq V\mathfrak{h}_{0}-C$. Furthermore, by Theorem 4.5 we can assume that the pocset $\mathcal{P}(\mathfrak{h}_{0},C)$ has no pairs of transverse elements.
4.3 The trees $T_{e}$
For brevity we will write $G_{0}=G_{\mathfrak{h}_{0}}$ for the $G$-stabilizer of $\mathfrak{h}_{0}$ and we will write $\delta$ for the edge boundary $\delta C$. As in Subsection 4.1, we can form the pocset $\mathcal{P}(\mathfrak{h}_{0},C)$. Now consider the subpocset
$\mathcal{P}_{0}:=\{gC,V\mathfrak{h}_{0}-gC\mid g\in G_{0}\}\subseteq\mathcal{P}(\mathfrak{h}_{0},C).$
Recall that the partial order on $\mathcal{P}_{0}$ is inclusion and the involution is complementation in $V\mathfrak{h}_{0}$, i.e. the map $D\mapsto V\mathfrak{h}_{0}-D$. We know that $\mathcal{P}(\mathfrak{h}_{0},C)$ has no pairs of transverse elements, so $\mathcal{P}_{0}$ also has no pairs of transverse elements. Therefore, the cubing of $\mathcal{P}_{0}$ is a tree, call it $T_{0}$.
Lemma 4.11.
$G_{0}$ acts on $T_{0}$ cocompactly and with finite edge stabilizers.
Proof.
The edges of $T_{0}$ correspond to the pairs $\{gC,V\mathfrak{h}_{0}-gC\}$ for $g\in G_{0}$, or equivalently to the edge boundaries $g\delta$ for $g\in G_{0}$. Hence $G_{0}$ acts on $T_{0}$ with a single orbit of edges and at most two orbits of vertices. Since $G_{0}$ acts properly on $\mathfrak{h}_{0}$, and since the sets of edges $g\delta$ are finite, we deduce that $G_{0}$ acts on $T_{0}$ with finite edge stabilizers. ∎
Lemma 4.12.
The action $G_{0}\curvearrowright T_{0}$ has no fixed point.
Proof.
It suffices to find $D\in\mathcal{P}_{0}$ and $g\in G_{0}$ with $gD\subsetneq D$, as then $g$ will have no fixed point in $T_{0}$. Indeed, $g$ will translate along the axis of $T_{0}$ that contains the edges corresponding to $\{g^{i}D,V\mathfrak{h}_{0}-g^{i}D\}$ for $i\in\mathbb{Z}$.
Recall that $C\cap V\hat{\mathfrak{h}}_{0}$ and $V\hat{\mathfrak{h}}_{0}-C$ are both unbounded by Lemma 4.7. As $G_{0}$ acts cocompactly on $\hat{\mathfrak{h}}_{0}$, and as the $G_{0}$-translates of $\delta$ all contain edges in $\hat{\mathfrak{h}}_{0}$ (Lemma 4.7 again), there exist $g_{1},g_{2}\in G_{0}$ with $g_{1}\delta$ contained in the induced subgraph of $C$ and $g_{2}\delta$ contained in the induced subgraph of $V\mathfrak{h}_{0}-C$. We deduce that one of the components of $\mathfrak{h}_{0}-g_{1}\delta$ contains both $\delta$ and $V\mathfrak{h}_{0}-C$, and the other component is contained in the induced subgraph of $C$. So either $g_{1}C\subsetneq C$ or $V\mathfrak{h}_{0}-C\subsetneq g_{1}C$. Similarly, we have $g_{2}(V\mathfrak{h}_{0}-C)\subsetneq V\mathfrak{h}_{0}-C$ or $C\subsetneq g_{2}(V\mathfrak{h}_{0}-C)$. If $g_{1}C\subsetneq C$ or $g_{2}(V\mathfrak{h}_{0}-C)\subsetneq V\mathfrak{h}_{0}-C$ then we are done, otherwise $V\mathfrak{h}_{0}-C\subsetneq g_{1}C$ and $C\subsetneq g_{2}(V\mathfrak{h}_{0}-C)$, so $g_{2}C\subsetneq V\mathfrak{h}_{0}-C\subsetneq g_{1}C$, and we are again done because $g_{1}^{-1}g_{2}C\subsetneq C$. ∎
If $\{g_{i}\mid i\in\Omega\}$ is a left transversal of $G_{0}$ in $G$ then we get an induced action of $G$ on the product $T_{0}\times G/G_{0}$, explicitly this is given by
$g\cdot(v,g_{i}G_{0}):=(g_{0}v,g_{j}G_{0}),$ (7)
where $gg_{i}=g_{j}g_{0}$ with $i,j\in\Omega$ and $g_{0}\in G_{0}$, and $g_{0}v$ refers to the action of $G_{0}$ on $T_{0}$. (This construction is essentially the same as the notion of induced representation from representation theory.) We may assume that the transversal $\{g_{i}\}$ includes the identity element, in which case the action of $G_{0}$ on $T_{0}\times\{G_{0}\}$ recovers the original action of $G_{0}$ on $T_{0}$.
We can then define the trees $T_{e}$ required for Theorem 4.1, and the action of $G$ on $\sqcup_{e}T_{e}$. Indeed if $e_{0}\in ET$ is the edge corresponding to $\mathfrak{h}_{0}$ (i.e. $p(\mathfrak{h}_{0})\subseteq T$ is bounded by $\hat{e}_{0}$), then put
$T_{g_{i}e_{0}}:=T_{0}\times\{g_{i}G_{0}\}$
for each $i\in\Omega$. Note that the $G$-stabilizer of $T_{e_{0}}$ is $G_{0}=G_{e_{0}}$, and the action $G_{0}\curvearrowright T_{e_{0}}$ is a non-trivial finite splitting with finite edge stabilizers by Lemmas 4.11 and 4.12. Similarly, the $G$-stabilizer of $T_{g_{i}e_{0}}$ is $g_{i}G_{0}g_{i}^{-1}=G_{g_{i}e_{0}}$, and the action $G_{g_{i}e_{0}}\curvearrowright T_{g_{i}e_{0}}$ is also a non-trivial finite splitting with finite edge stabilizers because it is conjugate to the action $G_{0}\curvearrowright T_{e_{0}}$. For $e\notin G\cdot e_{0}$ we define $T_{e}$ to be a single vertex and we define $gT_{e}=T_{ge}$. Altogether this gives us an action of $G$ on $\sqcup_{e}T_{e}$ that is compatible with the action on $T$, and that satisfies properties (a) and (b) from Theorem 4.1.
4.4 The pocset $\mathcal{P}$
Define an equivalence relation on $V\mathfrak{h}_{0}$ where $x\sim y$ if $x$ and $y$ are not separated in $\mathfrak{h}_{0}$ by any $g\delta$ with $g\in G_{0}$. Equivalently, $x\sim y$ if there is no $D\in\mathcal{P}_{0}$ with $x\in D$ and $y\notin D$. Let $[x]$ denote the equivalence class of $x$, and let $[x]^{*}=VX-[x]$.
Since $T_{0}$ is the cubing of $\mathcal{P}_{0}$, and $\mathcal{P}_{0}$ forms a wallspace $(V\mathfrak{h}_{0},\mathcal{P}_{0})$, we have a map $\lambda:V\mathfrak{h}_{0}\to VT_{0}$ as in Definition 3. Note that the equivalence classes $[x]$ are just the non-empty fibers of $\lambda$.
Lemma 4.13.
Each halfspace $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$ is contained in a $\sim$-class $[x]$.
Proof.
Fix $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$. Given $D\in\mathcal{P}_{0}$, we can write $D=gC$ or $D=V\mathfrak{h}_{0}-gC$ for some $g\in G_{0}$. Note that $g^{-1}\mathfrak{h}$ is another halfspace strictly contained in $\mathfrak{h}_{0}$, so our choice of $(\mathfrak{h}_{0},C)$ at the end of Subsection 4.2 implies that we either have $g^{-1}V\mathfrak{h}\subseteq C$ or $g^{-1}V\mathfrak{h}\subseteq V\mathfrak{h}_{0}-C$. Left multiplying by $g$ then implies that we either have $V\mathfrak{h}\subseteq D$ or $V\mathfrak{h}\subseteq V\mathfrak{h}_{0}-D$. This holds for all $D\in\mathcal{P}_{0}$, so the lemma follows. ∎
The subgraph of $X$ spanned by vertices that lie in $\mathfrak{h}_{0}-\hat{\mathfrak{h}}_{0}$ and not in halfspaces $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$ is the $p$-preimage of a vertex in $T$, so it is a connected subgraph (see Definition 2). As $\delta$ is contained in $\mathfrak{h}_{0}$ but avoids halfspaces $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$, there exists a finite connected subgraph $F\subseteq\mathfrak{h}_{0}$ that contains $\delta$ but avoids halfspaces $\mathfrak{h}\subsetneq\mathfrak{h}_{0}$. Let $L$ be the length of the longest embedded path in $F$.
Lemma 4.14.
Let $g\in G$ with $g\mathfrak{h}_{0}^{*}\subseteq\mathfrak{h}_{0}$. Let $[x]$ and $[y]$ be the $\sim$-classes with $g\mathfrak{h}_{0}^{*}\subseteq[x]$ and $\mathfrak{h}_{0}^{*}\subseteq g[y]$ (which exist by Lemma 4.13). Then $g[y]^{*}$ is contained in a bounded neighborhood of $[x]$.
Proof.
Observe that $g[y]^{*}$ is contained in the union of $g\mathfrak{h}_{0}^{*}$ and $g[z]$ for $\sim$-classes $[z]\neq[y]$. We already know that $g\mathfrak{h}_{0}^{*}\subseteq[x]$, so it remains to show that the sets $g[z]$ for $[z]\neq[y]$ are contained in a bounded neighborhood of $[x]$.
Let $z\in V\mathfrak{h}_{0}$ with $[z]\neq[y]$. Suppose $gz\notin[x]$. Suppose $g_{0}\delta$ separates $[z]$ from $[y]$ in $\mathfrak{h}_{0}$, with $g_{0}\in G_{0}$. So $gg_{0}\delta$ separates $gz$ from $g[y]$ in $g\mathfrak{h}_{0}$. Since $\mathfrak{h}_{0}^{*}\subseteq g[y]$, we have that $gz\in\mathfrak{h}_{0}$ and that $gg_{0}\delta$ separates $gz$ from $\mathfrak{h}_{0}^{*}$ in $g\mathfrak{h}_{0}$. As $gz\notin[x]$, there is some $g_{1}\in G_{0}$ such that $g_{1}\delta$ separates $gz$ from $[x]$ in $\mathfrak{h}_{0}$. And as $g\mathfrak{h}_{0}^{*}\subseteq[x]$, we know that $g_{1}\delta$ also separates $gz$ from $g\mathfrak{h}_{0}^{*}$ in $\mathfrak{h}_{0}$.
Let $\gamma$ be a path in $X$ from $gz$ to $[x]$. We claim that $\gamma$ intersects the $L$-neighborhood of $gg_{0}\delta$. Suppose not. As $gg_{0}\delta$ separates $gz$ from $\mathfrak{h}_{0}^{*}$ in $g\mathfrak{h}_{0}$, and $g_{1}\delta$ separates $gz$ from $g\mathfrak{h}_{0}^{*}$ in $\mathfrak{h}_{0}$, we see that $\gamma$ will stay in $\mathfrak{h}_{0}\cap g\mathfrak{h}_{0}$ unless it intersects $g_{1}\delta$. Since $g_{1}\delta$ separates $gz$ from $[x]$ in $\mathfrak{h}_{0}$, we deduce that $\gamma$ must intersect $g_{1}\delta$. Let $z^{\prime}$ be the first point of intersection. Observe that $gg_{0}\delta$ separates $z^{\prime}$ from $\mathfrak{h}_{0}^{*}$ in $g\mathfrak{h}_{0}$. We also have $z^{\prime}\in g_{1}F$, so there is a path $\gamma^{\prime}$ in $g_{1}F$ from $z^{\prime}$ to $\hat{\mathfrak{h}}_{0}\subseteq\mathfrak{h}_{0}^{*}$ of length at most $L$. We have $g_{1}F\subseteq g\mathfrak{h}_{0}$ by construction of $F$, so $\gamma^{\prime}$ must cross $gg_{0}\delta$. Hence $\gamma$ intersects the $L$-neighborhood of $gg_{0}\delta$, as claimed.
We have shown that either $d(gz,gg_{0}\delta)\leq L$ or the $L$-neighborhood of $gg_{0}\delta$ separates $gz$ from $[x]$ in $X$. Suppose for the moment we are in the latter case. As $X$ is one-ended, the $L$-neighborhood of $\delta$ only has one unbounded complementary component. Since $X$ is locally finite, there are only finitely many bounded complementary components, so there is a constant $K\geq L$ such that the bounded complementary components are contained in the $K$-neighborhood of $\delta$. Since $g\mathfrak{h}_{0}^{*}\subseteq[x]$, we know that $[x]$ is unbounded, so $gz$ is in a bounded complementary component of the $L$-neighborhood of $gg_{0}\delta$, and $d(gz,gg_{0}\delta)<K$. Thus in either case we get $d(gz,gg_{0}\delta)<K$. Then $d(gz,g\mathfrak{h}_{0}^{*})<K+L$, which in turn implies that $d(gz,[x])<K+L$. ∎
$gg_{0}\delta$$\mathfrak{h}_{0}$$g\mathfrak{h}_{0}$$g[y]$$gz$
$g_{1}\delta$$\mathfrak{h}_{0}$$g\mathfrak{h}_{0}$$[x]$$gz$$z^{\prime}$$\gamma$
Figure 2: Cartoon picture for the proof of Lemma 4.14.
We now build a pocset $\mathcal{P}$ using the halfspaces of $X$ and the $\sim$-classes. The partial order on $\mathcal{P}$ will be based on coarse inclusion of subsets of $X$. Specifically, for subsets $\mathfrak{a},\mathfrak{b}\subseteq X$, write $\mathfrak{a}\preceq\mathfrak{b}$ if $\mathfrak{a}$ is contained in a bounded neighborhood of $\mathfrak{b}$. Note that $\mathfrak{a}\subseteq\mathfrak{b}$ implies $\mathfrak{a}\preceq\mathfrak{b}$, and $\mathfrak{a}\preceq\mathfrak{b}\preceq\mathfrak{c}$ implies $\mathfrak{a}\preceq\mathfrak{c}$. By convention, $\emptyset\preceq\mathfrak{b}$ for any $\mathfrak{b}$.
Let
#=$\mathcal{H}^{\#}=\{\mathfrak{h}\in\mathcal{H}(X)\mid\hat{\mathfrak{h}}\notin G\cdot\hat{\mathfrak{h}}_{0}\},$
and let $\mathcal{M}_{0}$ denote the set of $\sim$-classes in $V\mathfrak{h}_{0}$ that are not contained in a bounded neighborhood of $\hat{\mathfrak{h}}_{0}$. Note that the $\sim$-class $[x]$ appearing in Lemma 4.13 is in $\mathcal{M}_{0}$, because the halfspace $\mathfrak{h}\subseteq[x]$ is not contained in a bounded neighborhood of $\hat{\mathfrak{h}}_{0}$. The same goes for the $\sim$-classes $[x]$ and $[y]$ appearing in Lemma 4.14. We also get the following lemma.
Lemma 4.15.
We cannot have $[x]\preceq[x]^{*}$ for $[x]\in\mathcal{M}_{0}$.
Proof.
Any path from $[x]$ to $[x]^{*}$ must either pass through $\hat{\mathfrak{h}}_{0}$ or through a $G_{0}$-translate of $\delta$, which is contained in a bounded neighborhood of $\hat{\mathfrak{h}}_{0}$. The lemma follows since $[x]$ is not contained in a bounded neighborhood of $\hat{\mathfrak{h}}_{0}$ (which is by definition of $\mathcal{M}_{0}$). ∎
Now define $\mathcal{P}$ to be the set of all pairs $(\mathfrak{a},\mathfrak{h})$, with $\mathfrak{a}\subseteq VX$ and $\mathfrak{h}\in\mathcal{H}(X)$, that arise in one of the following three ways:
$(\mathfrak{a},\mathfrak{h})=\begin{cases}(\mathfrak{h},\mathfrak{h}),&\mathfrak{h}\in\mathcal{H}^{\#},\\ (g[x],g\mathfrak{h}_{0}),&g\in G,\,[x]\in\mathcal{M}_{0},\\ (g[x]^{*},g\mathfrak{h}_{0}^{*}),&g\in G,\,[x]\in\mathcal{M}_{0},\end{cases}$ (8)
Define a relation $\leq$ on $\mathcal{P}$, where $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$ if
  1. (1)
    $\mathfrak{a}_{1}\prec\mathfrak{a}_{2}$ (meaning $\mathfrak{a}_{1}\preceq\mathfrak{a}_{2}$ but $\mathfrak{a}_{2}\npreceq\mathfrak{a}_{1}$), or
  2. (2)
    $\mathfrak{a}_{1}\preceq\mathfrak{a}_{2}$ and $\mathfrak{h}_{1}\subseteq\mathfrak{h}_{2}$.
Remark 4.16.
It can happen that conditions (1) and (2) both hold.
Lemma 4.17.
For distinct $[x],[y]\in\mathcal{M}_{0}$, we have $([x],\mathfrak{h}_{0})\leq([y]^{*},\mathfrak{h}_{0}^{*})$, but we do not have $([x],\mathfrak{h}_{0})\leq([y],\mathfrak{h}_{0})$ or $([x]^{*},\mathfrak{h}_{0}^{*})\leq([y]^{*},\mathfrak{h}_{0}^{*})$.
Proof.
We know that $[x]$ and $[y]$ are disjoint, so $[x]\preceq[y]^{*}$. Moreover, $\mathfrak{h}_{0}^{*}-\hat{\mathfrak{h}}_{0}\subseteq[y]^{*}$ is not contained in a bounded neighborhood of $[x]$, so $[x]\prec[y]^{*}$ and $([x],\mathfrak{h}_{0})\leq([y]^{*},\mathfrak{h}_{0}^{*}h)^{*}$ is an inequality of type (1).
Secondly, we cannot have $([x],\mathfrak{h}_{0})\leq([y],\mathfrak{h}_{0})$ as that would imply $[x]\preceq[y]\subseteq[x]^{*}$, contradicting Lemma 4.15. Similarly, we cannot have $([x]^{*},\mathfrak{h}_{0}^{*})\leq([y]^{*},\mathfrak{h}_{0}^{*})$ as that would imply $[y]\subseteq[x]^{*}\preceq[y]^{*}$, again contradicting Lemma 4.15. ∎
Lemma 4.18.
$\leq$ is a partial order on $\mathcal{P}$.
Proof.
Reflexivity holds by (2). Let’s now prove antisymmetry. If $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})\leq(\mathfrak{a}_{1},\mathfrak{h}_{1})$ then $\mathfrak{a}_{1}\preceq\mathfrak{a}_{2}\preceq\mathfrak{a}_{1}$, so both $\leq$ inequalities must be of type (2), which means that $\mathfrak{h}_{1}\subseteq\mathfrak{h}_{2}\subseteq\mathfrak{h}_{1}$, and so $\mathfrak{h}_{1}=\mathfrak{h}_{2}$. If $\mathfrak{h}_{1}\in\mathcal{H}^{\#}$ then $(\mathfrak{a}_{1},\mathfrak{h}_{1})=(\mathfrak{a}_{2},\mathfrak{h}_{2})=(\mathfrak{h}_{1},\mathfrak{h}_{1})$. If $\mathfrak{h}_{1}=g\mathfrak{h}_{0}$ (some $g\in G$), then $(\mathfrak{a}_{1},\mathfrak{h}_{1})=(g[x],g\mathfrak{h}_{0})$ and $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[y],g\mathfrak{h}_{0})$ for some $[x],[y]\in\mathcal{M}_{0}$, and $[x]=[y]$ by Lemma 4.17. Similarly, if $\mathfrak{h}_{1}=g\mathfrak{h}_{0}^{*}$ (some $g\in G$), then $(\mathfrak{a}_{1},\mathfrak{h}_{1})=(g[x]^{*},g\mathfrak{h}_{0}^{*})$ and $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[y]^{*},g\mathfrak{h}_{0}^{*})$ for some $[x],[y]\in\mathcal{M}_{0}$, and $[x]=[y]$ by Lemma 4.17.
Finally, we prove transitivity. Suppose $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})\leq(\mathfrak{a}_{3},\mathfrak{h}_{3})$. If one of the inequalities is of type (1), then $\mathfrak{a}_{1}\prec\mathfrak{a}_{3}$ and so $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{3},\mathfrak{h}_{3})$ is an inequality of type (1). On the other hand, if both inequalities are of type (2), then $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{3},\mathfrak{h}_{3})$ is also an inequality of type (2). ∎
Define an involution $(\mathfrak{a},\mathfrak{h})\mapsto(\mathfrak{a},\mathfrak{h})^{*}$ on $\mathcal{P}$ by
$(\mathfrak{h},\mathfrak{h})$ $\mapsto(\mathfrak{h}^{*},\mathfrak{h}^{*}),$ $\mathfrak{h}\in\mathcal{H}^{\#},$
$(g[x],g\mathfrak{h}_{0})$ $\mapsto(g[x]^{*},g\mathfrak{h}_{0}^{*}),$ $g\in G,\,[x]\in\mathcal{M}_{0},$
$(g[x]^{*},g\mathfrak{h}_{0}^{*})$ $\mapsto(g[x],g\mathfrak{h}_{0}),$ $g\in G,\,[x]\in\mathcal{M}_{0}.$
We have an action of $G$ on $\mathcal{P}$ given by $g\cdot(\mathfrak{a},\mathfrak{h}):=(g\mathfrak{a},g\mathfrak{h})$, which clearly respects $\leq$ and *. We will show that $(\mathcal{P},\leq,*)$ is a pocset with no pair of transverse elements. First we need the following lemma.
Lemma 4.19.
We cannot have both $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$ and $(\mathfrak{a}_{1},\mathfrak{h}_{1})^{*}\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$.
Proof.
If both inequalities did hold, then we would have $\mathfrak{a}_{1}\preceq\mathfrak{a}_{2}$ and $\mathfrak{a}_{1}^{*}\preceq\mathfrak{a}_{2}$, so $X\preceq\mathfrak{a}_{2}$ and $\mathfrak{a}_{2}^{*}\preceq\mathfrak{a}_{2}$. If $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ takes the form $(\mathfrak{h},\mathfrak{h})$ or $(g[x],g\mathfrak{h}_{0})$, then $\mathfrak{a}_{2}\subseteq\mathfrak{h}_{2}$ and we get a contradiction since $\mathfrak{h}_{2}^{*}\npreceq\mathfrak{h}_{2}$ (Lemma 3.2). If $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ takes the form $(g[x]^{*},g\mathfrak{h}_{0}^{*})$, then $\mathfrak{a}_{2}^{*}=g[x]\preceq g[x]^{*}$, contradicting Lemma 4.15. ∎
Lemma 4.20.
$(\mathcal{P},\leq,*)$ is a pocset with no pair of transverse elements.
Proof.
Firstly, by Lemma 4.19 and reflexivity of $\leq$ we can never have $(\mathfrak{a},\mathfrak{h})\leq(\mathfrak{a},\mathfrak{h})^{*}$. It remains to show that $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$ implies $(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}\leq(\mathfrak{a}_{1},\mathfrak{h}_{1})^{*}$, and that $\mathcal{P}$ has no pair of transverse elements. Equivalently, we must show that for all distinct $(\mathfrak{a}_{1},\mathfrak{h}_{1}),(\mathfrak{a}_{2},\mathfrak{h}_{2})\in\mathcal{P}$ exactly two of the following inequalities hold, and that these two inequalities are in the same row.
$(\mathfrak{a}_{1},\mathfrak{h}_{1})$ $\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$ $(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}$ $\leq(\mathfrak{a}_{1},\mathfrak{h}_{1})^{*}$ (9)
$(\mathfrak{a}_{1},\mathfrak{h}_{1})$ $\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}$ $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ $\leq(\mathfrak{a}_{1},\mathfrak{h}_{1})^{*}$
$(\mathfrak{a}_{1},\mathfrak{h}_{1})^{*}$ $\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$ $(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}$ $\leq(\mathfrak{a}_{1},\mathfrak{h}_{1})$
$(\mathfrak{a}_{1},\mathfrak{h}_{1})^{*}$ $\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}$ $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ $\leq(\mathfrak{a}_{1},\mathfrak{h}_{1})$
To begin, let’s show that if one of the rows of (9) holds then all inequalities in the other rows fail. Without loss of generality, assume that the first row of (9) holds. Then $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}$ fails, else we get $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}\leq(\mathfrak{a}_{1},\mathfrak{h}_{1})^{*}$, contradicting what we said at the start of the proof. The other inequalities in the second and third rows of (9) fail similarly. And the inequalities in the fourth row fail because otherwise we could combine them with the inequalities in the first row and deduce that $(\mathfrak{a}_{1},\mathfrak{h}_{1})=(\mathfrak{a}_{2},\mathfrak{h}_{2})$ using antisymmetry of $\leq$.
We will spend the rest of the proof showing that one of the rows of (9) holds. Among elements of the form $(\mathfrak{h},\mathfrak{h})$ in $\mathcal{P}$ (with $\mathfrak{h}\in\mathcal{H}^{\#}$), the relation $\leq$ is determined by the projection of these elements to the pocset $(\mathcal{H}(X),\subseteq)$. Since this pocset has no pairs of transverse elements, we deduce that if $(\mathfrak{a}_{1},\mathfrak{h}_{1}),(\mathfrak{a}_{2},\mathfrak{h}_{2})$ are both of the form $(\mathfrak{h},\mathfrak{h})$ with $\mathfrak{h}\in\mathcal{H}^{\#}$, then one of the rows of (9) holds.
Henceforth, assume that $(\mathfrak{a}_{1},\mathfrak{h}_{1}),(\mathfrak{a}_{2},\mathfrak{h}_{2})$ are not both of the form $(\mathfrak{h},\mathfrak{h})$ with $\mathfrak{h}\in\mathcal{H}^{\#}$. Without loss of generality, assume $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ is not of this form, so $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ is of the form $(g[x],g\mathfrak{h}_{0})$ or $(g[x]^{*},g\mathfrak{h}_{0}^{*})$. Since $G$ acts on $\mathcal{P}$ respecting $\leq$ and *, it suffices to consider the case where $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ is of the form $([x],\mathfrak{h}_{0})$ or $([x]^{*},\mathfrak{h}_{0}^{*})$. And replacing $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ with $(\mathfrak{a}_{1},\mathfrak{h}_{1})^{*}$ if necessary, we may assume $(\mathfrak{a}_{1},\mathfrak{h}_{1})=([x],\mathfrak{h}_{0})$. We now have three cases to consider:
  1. (1)
    Suppose $\mathfrak{h}_{2}\in\{\mathfrak{h}_{0},\mathfrak{h}_{0}^{*}\}$. Replacing $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ with $(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}$ if necessary, we may assume that $\mathfrak{h}_{2}=\mathfrak{h}_{0}$. Then $(\mathfrak{a}_{2},\mathfrak{h}_{2})=([y],\mathfrak{h}_{0})$ for some $[y]\in\mathcal{M}_{0}$. Since $(\mathfrak{a}_{1},\mathfrak{h}_{1}),(\mathfrak{a}_{2},\mathfrak{h}_{2})$ are distinct, we have $[x]\neq[y]$, so the second row of (9) holds by Lemma 4.17.
  2. (2)
    Suppose $\mathfrak{h}_{2}\subsetneq\mathfrak{h}_{0}$ or $\mathfrak{h}_{2}^{*}\subsetneq\mathfrak{h}_{0}$. Replacing $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ with $(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}$ if necessary, we may assume that $\mathfrak{h}_{2}\subsetneq\mathfrak{h}_{0}$. By Lemma 4.13, there is a unique $[y]\in\mathcal{M}_{0}$ with $\mathfrak{h}_{2}\subseteq[y]$. Let’s first show that either the third or fourth row of (9) holds in the case $[x]=[y]$. We split into three subcases:
    1. (a)
      Suppose $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(\mathfrak{h}_{2},\mathfrak{h}_{2})$. Then $\mathfrak{a}_{2}=\mathfrak{h}_{2}\subseteq[y]=\mathfrak{a}_{1}$, so the fourth row of (9) holds (and they are inequalities of type (2)).
    2. (b)
      Suppose $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[z],g\mathfrak{h}_{0})$, with $g\in G$, $[z]\in\mathcal{M}_{0}$. Then $\mathfrak{a}_{2}=g[z]\subseteq\mathfrak{h}_{2}\subseteq[y]=\mathfrak{a}_{1}$, so again the fourth row of (9) holds.
    3. (c)
      Suppose $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[z]^{*},g\mathfrak{h}_{0}^{*})$, with $g\in G$, $[z]\in\mathcal{M}_{0}$. Let $[y^{\prime}]$ be the $\sim$-class with $\mathfrak{h}_{0}^{*}\subseteq g[y^{\prime}]$. By Lemma 4.14 (applied both ways round), we have $g[y^{\prime}]^{*}\preceq[y]$ and $[y]^{*}\preceq g[y^{\prime}]$, so
      $(g[y^{\prime}]^{*},g\mathfrak{h}_{0}^{*})\leq(\mathfrak{a}_{1},\mathfrak{h}_{1})\quad\text{and}\quad(\mathfrak{a}_{1},\mathfrak{h}_{1})^{*}\leq(g[y^{\prime}],g\mathfrak{h}_{0}).$ (10)
      If $[z]=[y^{\prime}]$, then the fourth row of (9) holds by (10). If $[z]\neq[y^{\prime}]$, then
      $(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}\leq(g[y^{\prime}]^{*},g\mathfrak{h}_{0}^{*})\quad\text{and}\quad(g[y^{\prime}],g\mathfrak{h}_{0})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$ (11)
      by Lemma 4.17, and the third row of (9) holds by combining (10) and (11).
    Now suppose $[x]\neq[y]$. Then
    $(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq([y],\mathfrak{h}_{0})^{*}\quad\text{and}\quad([y],\mathfrak{h}_{0})\leq(\mathfrak{a}_{1},\mathfrak{h}_{1})^{*}$ (12)
    by Lemma 4.17. By what we showed above for the case $[x]=[y]$, we know that either the third or fourth row of (9) holds for the pair $([y],\mathfrak{h}_{0}),(\mathfrak{a}_{2},\mathfrak{h}_{2})$, namely we must be in one of the following two subcases:
    1. (i)
      Suppose $([y],\mathfrak{h}_{0})^{*}\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$ and $(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}\leq([y],\mathfrak{h}_{0})$. Combining these inequalities with (12) we deduce that the first row of (9) holds for $(\mathfrak{a}_{1},\mathfrak{h}_{1}),(\mathfrak{a}_{2},\mathfrak{h}_{2})$.
    2. (ii)
      Suppose $([y],\mathfrak{h}_{0})^{*}\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}$ and $(\mathfrak{a}_{2},\mathfrak{h}_{2})\leq([y],\mathfrak{h}_{0})$. Combining these inequalities with (12) we deduce that the second row of (9) holds for $(\mathfrak{a}_{1},\mathfrak{h}_{1}),(\mathfrak{a}_{2},\mathfrak{h}_{2})$.
  3. (3)
    Suppose $\mathfrak{h}_{0}\subsetneq\mathfrak{h}_{2}$ or $\mathfrak{h}_{0}\subsetneq\mathfrak{h}_{2}^{*}$. Replacing $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ with $(\mathfrak{a}_{2},\mathfrak{h}_{2})^{*}$ if necessary, we may assume that $\mathfrak{h}_{0}\subsetneq\mathfrak{h}_{2}$. Again, we split into three subcases:
    1. (a)
      If $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(\mathfrak{h}_{2},\mathfrak{h}_{2})$, then $[x]\subseteq\mathfrak{h}_{0}\subseteq\mathfrak{h}_{2}$, so the first row of (9) holds (and they are both inequalities of type (2)).
    2. (b)
      If $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[z],g\mathfrak{h}_{0})$, with $g\in G$ and $[z]\in\mathcal{M}_{0}$, then up to the action of $G$ (and with the roles of $(\mathfrak{a}_{1},\mathfrak{h}_{1}),(\mathfrak{a}_{2},\mathfrak{h}_{2})$ reversed) this is the same as case (2).
    3. (c)
      Finally, if $(\mathfrak{a}_{2},\mathfrak{h}_{2})=(g[z]^{*},g\mathfrak{h}_{0}^{*})$, with $g\in G$ and $[z]\in\mathcal{M}_{0}$, then $[x]\subseteq\mathfrak{h}_{0}\subseteq\mathfrak{h}_{2}\preceq g[z]^{*}$, so the first row of (9) holds (and they are both inequalities of type (2)).∎
4.5 The tree $T^{\prime}$
Let $T^{\prime}$ be the cubing of $\mathcal{P}$. Lemma 4.20 tells us that $T^{\prime}$ is a tree. Moreover, the action of $G$ on $\mathcal{P}$ induces an action of $G$ on $T^{\prime}$. The edges in $T^{\prime}$ correspond to pairs $\{(\mathfrak{a},\mathfrak{h}),(\mathfrak{a},\mathfrak{h})^{*}\}\subseteq\mathcal{P}$, and, since $(\mathfrak{a},\mathfrak{h})^{*}$ always takes the form $(\mathfrak{a}^{\prime},\mathfrak{h}^{*})$, an edge inversion in $T^{\prime}$ would mean there is an element $g\in G$ with $g\mathfrak{h}=\mathfrak{h}^{*}$. But that would give rise to an edge inversion in $T$, which doesn’t happen, hence $G$ acts on $T^{\prime}$ without edge inversions, i.e. $G\curvearrowright T^{\prime}$ is a splitting. We may assume that the action $G\curvearrowright T^{\prime}$ is minimal since we can always replace $T^{\prime}$ with a minimal $G$-invariant subtree. The description of edges in $T^{\prime}$ also makes it clear that every edge stabilizer of $T^{\prime}$ is contained in an edge stabilizer of $T$.
Lemma 4.21.
Every vertex stabilizer of $T^{\prime}$ is contained in a vertex stabilizer of $T$.
Proof.
We will prove this by constructing a $G$-equivariant map $\phi:VT^{\prime}\to VT$. As in Proposition 3.1, we will view the vertices of $T$ as DCC ultrafilters on $\mathcal{H}(X)$ (which is isomorphic to the pocset $\mathcal{H}(T)$) and the vertices of $T^{\prime}$ as DCC ultrafilters on $\mathcal{P}$. Given $v\in VT^{\prime}$, we then define
$\phi(v):=$$\{\mathfrak{h}\in\mathcal{H}^{\#}\mid(\mathfrak{h},\mathfrak{h})\in v\}$
$\cup\{g\mathfrak{h}_{0}\mid g\in G,\,(g[x],g\mathfrak{h}_{0})\in v\text{ for some }[x]\in\mathcal{M}_{0}\}$
$\cup\{g\mathfrak{h}_{0}^{*}\mid g\in G,\,(g[x]^{*},g\mathfrak{h}_{0}^{*})\in v\text{ for all }[x]\in\mathcal{M}_{0}\}.$
Note that the properties required for $g\mathfrak{h}_{0}$ and $g\mathfrak{h}_{0}^{*}$ to be in $\phi(v)$ only depend on the halfspaces $g\mathfrak{h}_{0}$ and $g\mathfrak{h}_{0}^{*}$, not on the element $g\in G$. The map $\phi:VT^{\prime}\to VT$ is $G$-equivariant by definition of the $G$-action on $\mathcal{P}$. Following Definition 3, we must show that $\phi(v)$ is complete, consistent, and DCC.
Let’s first prove completeness of $\phi(v)$. For $\mathfrak{h}\in\mathcal{H}^{\#}$, exactly one of $(\mathfrak{h},\mathfrak{h})$ and $(\mathfrak{h}^{*},\mathfrak{h}^{*})$ is in $v$ by completeness of $v$, so exactly one of $\mathfrak{h}$ and $\mathfrak{h}^{*}$ is in $\phi(v)$. For $g\in G$ and $[x]\in\mathcal{M}_{0}$, exactly one of $(g[x],g\mathfrak{h}_{0})$ and $(g[x]^{*},g\mathfrak{h}_{0}^{*})$ is in $v$ by completeness of $v$, so for each $g\in G$ we have exactly one of $g\mathfrak{h}_{0}$ and $g\mathfrak{h}_{0}^{*}$ in $\phi(v)$.
Before showing that $\phi(v)$ is consistent and DCC, we make the following observation, which follows from Lemmas 4.13 and 4.14: if $\mathfrak{h}_{1}\subsetneq\mathfrak{h}_{2}$ in $\mathcal{H}(X)$ then
$(\mathfrak{a}_{1},\mathfrak{h}_{1})\leq(\mathfrak{a}_{2},\mathfrak{h}_{2})$ (13)
in $\mathcal{P}$, where
  1. (1)
    if $\mathfrak{h}_{1}\in\mathcal{H}^{\#}$ then $\mathfrak{a}_{1}=\mathfrak{h}_{1}$,
  2. (2)
    if $\mathfrak{h}_{1}=g_{1}\mathfrak{h}_{0}$ then $\mathfrak{a}_{1}=g_{1}[x]$ and (13) holds for all choices of $[x]\in\mathcal{M}_{0}$,
  3. (3)
    if $\mathfrak{h}_{1}=g_{1}\mathfrak{h}_{0}^{*}$ then $\mathfrak{a}_{1}=g_{1}[x]^{*}$ for the unique $[x]\in\mathcal{M}_{0}$ with $\mathfrak{h}_{2}^{*}\subseteq g_{1}[x]$,
  4. (4)
    if $\mathfrak{h}_{2}\in\mathcal{H}^{\#}$ then $\mathfrak{a}_{2}=\mathfrak{h}_{2}$,
  5. (5)
    if $\mathfrak{h}_{2}=g_{2}\mathfrak{h}_{0}$ then $\mathfrak{a}_{2}=g_{2}[y]$ for the unique $[y]\in\mathcal{M}_{0}$ with $\mathfrak{h}_{1}\subseteq g_{2}[y]$,
  6. (6)
    if $\mathfrak{h}_{2}=g_{2}\mathfrak{h}_{0}^{*}$ then $\mathfrak{a}_{2}=g_{2}[y]^{*}$ and (13) holds for all choices of $[y]\in\mathcal{M}_{0}$.
We now prove consistency of $\phi(v)$. If $\mathfrak{h}_{1}\subsetneq\mathfrak{h}_{2}$ and $\mathfrak{h}_{1}\in\phi(v)$ then $(\mathfrak{a}_{1},\mathfrak{h}_{1})\in v$, where $(\mathfrak{a}_{1},\mathfrak{h}_{1})$ is given by the appropriate case from (1)(3) (and in case (2) we only have $(g_{1}[x],g_{1}\mathfrak{h}_{0})\in v$ for one $[x]\in\mathcal{M}_{0}$). Consistency of $v$ then implies that $(\mathfrak{a}_{2},\mathfrak{h}_{2})\in v$, where $(\mathfrak{a}_{2},\mathfrak{h}_{2})$ is given by the appropriate case from (4)(6) (and in case (6) we have $(g_{2}[y]^{*},g_{2}\mathfrak{h}_{0}^{*})\in v$ for all $[y]\in\mathcal{M}_{0}$). It follows that $\mathfrak{h}_{2}\in\phi(v)$.
Finally, we prove that $\phi(v)$ is DCC. Suppose for contradiction that we have an infinite descending chain $\mathfrak{h}_{1}\supsetneq\mathfrak{h}_{2}\supsetneq\mathfrak{h}_{3}\supsetneq\dots$ in $\phi(v)$. Then we can turn this into an infinite descending chain
$(\mathfrak{a}_{1},\mathfrak{h}_{1})\geq(\mathfrak{a}_{2},\mathfrak{h}_{2})\geq(\mathfrak{a}_{3},\mathfrak{h}_{3})\geq\dots$ (14)
in $v$, where
  1. (1)
    if $\mathfrak{h}_{i}\in\mathcal{H}^{\#}$ then $\mathfrak{a}_{i}=\mathfrak{h}_{i}$,
  2. (2)
    if $\mathfrak{h}_{i}=g\mathfrak{h}_{0}$ then $\mathfrak{a}_{i}=g[x]$ for the unique $[x]\in\mathcal{M}_{0}$ with $\mathfrak{h}_{i+1}\subseteq g[x]$,
  3. (3)
    if $\mathfrak{h}_{i}=g\mathfrak{h}_{0}^{*}$ then $\mathfrak{a}_{i}=g[x]^{*}$ for the unique $[x]\in\mathcal{M}_{0}$ with $\mathfrak{h}_{i-1}^{*}\subseteq g[x]$ (or any $[x]\in\mathcal{M}_{0}$ if $i=1$).
(14) is a chain for the same reason that the inequality (13) holds. But of course this contradicts the fact that $v$ is DCC. ∎
Property (c) in Theorem 4.1 is given by the following lemma.
Lemma 4.22.
There is an injective $G$-equivariant map $\tau:ET^{\prime}\to\sqcup_{e}VT_{e}$.
Proof.
Recall that we have a wallspace $(V\mathfrak{h}_{0},\mathcal{P}_{0})$ with cubing $T_{0}$, and an associated map $\lambda:V\mathfrak{h}_{0}\to VT_{0}$. Moreover, it follows from the constructions that $\lambda$ is $G_{0}$-equivariant. We observed that the $\sim$-classes $[x]$ are precisely the non-empty fibers of $\lambda$, so in particular we get a well-defined injective $G_{0}$-equivariant map $\mathcal{M}_{0}\to VT_{0}$ given by $[x]\mapsto\lambda(x)$.
Now recall the construction of the trees $T_{e}$ ($e\in ET$) in Subsection 4.3, and the action of $G$ on $\sqcup_{e}T_{e}$. With $e_{0}$ being the edge corresponding to the halfspace $\mathfrak{h}_{0}$, the tree $T_{e_{0}}$ is identified with the tree $T_{0}$. We now extend the map $\mathcal{M}_{0}\to VT_{0}$ constructed above to a map $\tau:ET^{\prime}\to\sqcup_{e}VT_{e}$. We identify the edges of $T^{\prime}$ with pairs $\{(\mathfrak{a},\mathfrak{h}),(\mathfrak{a},\mathfrak{h})^{*}\}\subseteq\mathcal{P}$, so we have two types of edges to consider.
  • For $\mathfrak{h}\in\mathcal{H}^{\#}$, we have an edge $e^{\prime}=\{(\mathfrak{h},\mathfrak{h}),(\mathfrak{h}^{*},\mathfrak{h}^{*})\}\in ET^{\prime}$ and also an edge $e\in ET$. In this case $e\notin G\cdot e_{0}$, so the tree $T_{e}$ is just a single vertex, and we define $\tau(e^{\prime})$ to be this vertex. The maps #$\mathcal{H}^{\#}\to ET$ and #$\mathcal{H}^{\#}\to ET^{\prime}$ are $G$-equivariant, so $\tau(ge^{\prime})=g\cdot\tau(e^{\prime})$ for all $g\in G$.
  • Recall that we have a left transversal $\{g_{i}\mid i\in\Omega\}$ of $G_{0}$ in $G$, so the second type of edge in $T^{\prime}$ can be written in the form
    $e^{\prime}=\{(g_{i}[x],g_{i}\mathfrak{h}_{0}),(g_{i}[x]^{*},g_{i}\mathfrak{h}_{0}^{*})\},$
    with $i\in\Omega$ and $[x]\in\mathcal{M}_{0}$. We then define
    $\tau(e^{\prime}):=(\lambda(x),g_{i}G_{0})\in T_{0}\times\{g_{i}G_{0}\}=:T_{g_{i}e_{0}}.$
    To check $G$-equivariance, take $g\in G$, and let $gg_{i}=g_{j}g_{0}$ with $j\in\Omega$ and $g_{0}\in G_{0}$. The definition of $G\curvearrowright\sqcup_{e}T_{e}$ tells us that
    $g\cdot(\lambda(x),g_{i}G_{0})=(g_{0}\lambda(x),g_{j}G_{0}).$
    Meanwhile,
    $\tau(ge^{\prime})$ $=\tau(\{(gg_{i}[x],gg_{i}\mathfrak{h}_{0}),(gg_{i}[x]^{*},gg_{i}\mathfrak{h}_{0}^{*})\})$
    $=\tau(\{(g_{j}g_{0}[x],g_{j}g_{0}\mathfrak{h}_{0}),(g_{j}g_{0}[x]^{*},g_{j}g_{0}\mathfrak{h}_{0}^{*})\})$
    $=\tau(\{(g_{j}[g_{0}x],g_{j}\mathfrak{h}_{0}),(g_{j}[g_{0}x]^{*},g_{j}\mathfrak{h}_{0}^{*})\})$
    $=(\lambda(g_{0}x),g_{j}G_{0})$
    $=(g_{0}\lambda(x),g_{j}G_{0}).$
    So $\tau(ge^{\prime})=g\cdot\tau(e^{\prime})$, as required.
Finally, the map $\tau:ET^{\prime}\to\sqcup_{e}VT_{e}$ is injective because the map $\mathcal{M}_{0}\to VT_{0}$ given by $[x]\mapsto\lambda(x)$ is injective. ∎
It follows from Lemma 4.22 and property (a) of Theorem 4.1 that each edge stabilizer of $G\curvearrowright T^{\prime}$ is equal to a vertex stabilizer from one of the splittings $G_{e}\curvearrowright T_{e}$. In particular, the edge stabilizers of $G\curvearrowright T^{\prime}$ are finitely generated by Lemma 2.2. Finally, property (d) of Theorem 4.1 is given by the following lemma.
Lemma 4.23.
If no vertex stabilizer of $T$ fixes an edge in $T$, then there is a vertex $v_{0}\in VT$ such that $G_{v_{0}}$ has no fixed point in $T^{\prime}$.
Proof.
Recall that $e_{0}\in ET$ is the edge corresponding to the halfspace $\mathfrak{h}_{0}$, in particular $G_{0}=G_{e_{0}}$. The vertex $v_{0}\in VT$ will be the endpoint of $e_{0}$ that is contained in the halfspace $p(\mathfrak{h}_{0})\in\mathcal{H}(T)$. The vertex stabilizer $G_{v_{0}}$ fixes no edges in $T$, so there exists $g\in G_{v_{0}}-G_{e_{0}}$. Then $g\mathfrak{h}_{0}^{*}\subsetneq\mathfrak{h}_{0}$. By Lemma 4.14, there exist $[x],[y]\in\mathcal{M}_{0}$ with $g[y]^{*}\preceq[x]$, so $(g[y]^{*},g\mathfrak{h}_{0}^{*})\leq([x],\mathfrak{h}_{0})$ in $\mathcal{P}$. Recall that we have a $G_{0}$-equivariant map $\mathcal{M}_{0}\to T_{0}$ (as in the proof of Lemma 4.22), and $G_{0}$ has no fixed point in $T_{0}$ by Lemma 4.12, so the $G_{0}$-orbits in $\mathcal{M}_{0}$ are all infinite. In particular, there exists $g_{0}\in G_{0}$ with $g_{0}[x]\neq[y]$. Then $(g_{0}[x],\mathfrak{h}_{0})\leq([y]^{*},\mathfrak{h}_{0}^{*})$ by Lemma 4.17. Putting this all together, we get
$gg_{0}\cdot([x],\mathfrak{h}_{0})=(gg_{0}[x],g\mathfrak{h}_{0})\leq(g[y]^{*},g\mathfrak{h}_{0}^{*})\leq([x],\mathfrak{h}_{0}).$
Therefore $gg_{0}\in G_{v_{0}}$ acts hyperbolically on $T^{\prime}$, translating along the axis that contains the edges corresponding to $\{(gg_{0})^{i}\cdot([x],\mathfrak{h}_{0}),(gg_{0})^{i}\cdot([x]^{*},\mathfrak{h}_{0}^{*})\}$ for $i\in\mathbb{Z}$. As a result, $G_{v_{0}}$ has no fixed point in $T^{\prime}$. ∎
5 Groups $G$ that are simply connected at infinity; have $H^{2}(G,\mathbb{Z}G)=\{0\}$
We begin this section with basic definitions. The notion of simple connectivity at infinity for groups and spaces is a classical and well studied notion. We use [Geo08] as a basic reference.
Definition 13. The space $X$ is simply connected at infinity if for any compact set $C$ there is a compact set $D$ such that loops $\alpha:[0,1]\to X-D$ are homotopically trivial in $X-C$. This means there is a homotopy $H:[0,1]\times[0,1]\to X-C$ such that $H(t,0)=\alpha(t)$, $H(0,t)=H(1,t)=H(t,1)=H(0,0)$ for all $t\in[0,1]$.
Definition 14. The space $X$ is 1-acyclic at infinity (has pro-finite first homology at infinity) if for any compact set $C$ there is a compact set $D$ such that the image of $H_{1}(X-D)$ in $H_{1}(X-C)$ under the homomorphism induced by the inclusion of $X-D$ into $X-C$ is trivial (respectively, finite). A loop $\alpha$ in $X-C$ is homologically trivial if there is an orientable 2-manifold $M$ bounded by an embedded loop $\alpha^{\prime}$ and a continuous map $H:M\to X-C$ such that $H$ restricted to $\alpha^{\prime}$ maps (in the obvious way) to $\alpha$ (see Proposition 12.8, [GH81]).
Definition 15. A (proper) cellular map $f:X\to Y$ between CW-complexes is a (proper) $k$-equivalence if there is a (proper) cellular map $g:Y\to X$ such that $gf:X^{k-1}\to X$ is (properly) homotopic to the inclusion $X^{k-1}\to X$ and $fg:Y^{k-1}\to Y$ is (properly) homotopic to the inclusion $Y^{k-1}\to Y$.
Theorem 5.1 (See the first paragraph of §16.5,[Geo08]).
Suppose $X$ and $Y$ are finite connected CW-complexes with $\pi_{1}(X)$ isomorphic to $\pi_{1}(Y)$. Then there is a 2-equivalence between $X$ and $Y$, and hence (by lifting) a proper 2-equivalence between the 2-skeletons of the universal covers of $X$ and $Y$.
Theorem 5.2 (Proposition 16.2.3 and 16.1.14, [Geo08]).
If $X$ and $Y$ are locally finite, connected CW-complexes and there is a proper 2-equivalence $f:X\to Y$, then $X$ is simply connected at infinity, 1-acyclic at infinity, or has pro-finite first homology at infinity if and only if $Y$ does as well.
Definition 16. If $G$ is a finitely presented group and $X$ is some (equivalently any) connected finite complex with $\pi_{1}(X)$ isomorphic to $G$, then $G$ is simply connected at infinity, is 1-acyclic at infinity, or has pro-finite first homology at infinity if the same holds for the universal cover of $X$.
The next result follows directly from parts (ii) and (iii) of the main corollary of [GM86] in the special case with $n=2$. In particular, if the finitely presented group $G$ is simply connected at infinity, then it is 1-acyclic at infinity and if $G$ is 1-acyclic at infinity, then the first homology at infinity of $G$ is pro-finite.
Theorem 5.3 ([GM86]).
If $G$ is a finitely presented group, then $H^{2}(G,\mathbb{Z}G)=\{0\}$ if and only if the first homology at infinity of $G$ is pro-finite.
Finally we mention the following:
Theorem 5.4 (Theorem 18.2.11, [Geo08]).
If $G$ and $H$ are finitely presented quasi-isometric groups then $G$ is simply connected at infinity, is 1-acyclic at infinity, or has pro-finite first homology at infinity if and only if $H$ does as well.
6 Proof of Theorem 1.4
Theorem 1.4.
Suppose $X$ is a locally finite CW-complex, and $X_{1}$ and $X_{2}$ are connected one-ended subcomplexes of $X$ such that $X_{1}\cup X_{2}=X$. If $K$ is a finite subcomplex of $X$ (possibly empty) such that $(X_{1}\cap X_{2})-K$ has more than one unbounded component, then $X$ does not have pro-finite first homology at infinity. In particular, $X$ is not simply connected at infinity.
Proof.
Assume that $X$ has pro-finite first homology at infinity. Let $X_{0}=X_{1}\cap X_{2}$. Let $D$ be a finite subcomplex of $X$ containing $K$ such that for any loop $\tau$ in $X-D$, there is $n>0$ such that $\tau^{n}$ is homologically trivial in $X-K$. Since both $X_{1}$ and $X_{2}$ are one-ended, there is a finite subcomplex $E$ of $X$ containing $D$ such that any two vertices of $X_{0}-E$ can be joined by an edge path in $X_{1}-D$ and by an edge path in $X_{2}-D$. Choose vertices $v$ and $w$ in $X_{0}-E$ such that $v$ and $w$ cannot be joined by an edge path in $X_{0}-K$ (so $v\in V$ and $w\in W$ where $V$ and $W$ are distinct components of $X_{0}-K$). Let $\alpha$ be an edge path in $X_{1}-D$ from $v$ to $w$ and let $\beta$ be an edge path in $X_{2}-D$ from $w$ to $v$. The concatenation, $(\alpha,\beta)$, is a loop in $X-D$. By our earlier assumption, $(\alpha,\beta)^{n}$ is homologically trivial in $X-K$. Let $M$ be an orientable 2-manifold bounded by an embedded loop $(\alpha^{\prime}_{1},\beta_{1}^{\prime},\ldots,\alpha_{n}^{\prime},\beta_{n}^{\prime})$ and let $H:M\to X-K$ be a continuous map such that $H$ restricted to $\alpha_{i}^{\prime}$ and $\beta_{i}^{\prime}$ maps (in the obvious way) to the edge paths $\alpha$ and $\beta$ respectively. Let $v_{i}^{\prime}$ and $w_{i}^{\prime}$ be the initial and terminal points of $\alpha_{i}^{\prime}$ (See Figure 3).
Figure 3: The Relevant Maps
Lemma 6.1.
The sets $H^{-1}(W)$ and $H^{-1}(X_{0}-W)$ are disjoint closed subsets of $M$ containing $\{w_{1}^{\prime},\ldots w_{m}^{\prime}\}^{\prime}$ and $\{v_{1}^{\prime},\ldots,v_{n}^{\prime}\}$ respectively.
Proof.
First note that $H^{-1}(X_{0})$ is a closed subset of $M$. As $X_{0}$ is locally connected, each component of $X_{0}-K$ is open in $X_{0}$. In particular, $X_{0}-W$ is closed in $X_{0}$ and hence closed in $X$. Then $H^{-1}(X_{0}-W)$ is closed in $M$. The set $W\cup K$ is closed in $X_{0}$ and in $X$. As the image of $H$ misses $K$, $H^{-1}(W)=H^{-1}(W\cup K)$ and is closed in $M$. ∎
Consider the circle $S^{1}$ with diametrically opposite vertices $\bar{v}$ and $\bar{w}$. Let $T$ and $B$ be the two closed arcs in $S^{1}$ such that $T\cup B=S^{1}$ and $T\cap B=\{\bar{v},\bar{w}\}$. Map the (disjoint) closed sets $H^{-1}(W)$ and $H^{-1}(X_{0}-W)$ in $H^{-1}(X_{1})$ to $\{\bar{w}\}$ and $\{\bar{v}\}$ respectively. Extend this map by Tietze’s extension theorem to a continuous map $f_{1}:H^{-1}(X_{1})\to T$. Similarly define $f_{2}:H^{-1}(X_{2})\to B$ (with $f_{2}(H^{-1}(W))=\{\bar{w}\}$ and $f_{2}(H^{-1}(X_{0}-W))=\{\bar{v}\}$). Since $f_{1}$ and $f_{2}$ agree on the closed set $H^{-1}(X_{1})\cap H^{-1}(X_{2})=H^{-1}(X_{0})=H^{-1}(W)\cup H^{-1}(X_{0}-W)$, we have a continuous function $f:M\to S^{1}$ that agrees with $f_{1}$ and $f_{2}$ on $H^{-1}(X_{1})$ and $H^{-1}(X_{2})$ respectively. Note that for each $i$, $f\alpha_{i}^{\prime}=f_{1}\alpha_{i}^{\prime}=\bar{\alpha}_{i}$ is a path in $T$ from $\bar{v}$ to $\bar{w}$ and $f\beta_{i}^{\prime}=f_{2}\beta_{i}^{\prime}=\bar{\beta}_{i}$ is a path in $B$ from $\bar{w}$ to $\bar{v}$. Furthermore, for each $i$, the homotopy class of loop $(\bar{\alpha}_{i},\bar{\beta}_{i})$ at $\bar{v}$ generates $\pi_{1}(S_{1},\bar{v})=\mathbb{Z}$. The contradiction arises since the homotopy class of the loop $(\alpha_{1}^{\prime},\beta_{1}^{\prime},\ldots,\alpha_{n}^{\prime},\beta_{n}^{\prime})$ is a commutator in $\pi_{1}(M,v^{\prime})$ and so maps to the trivial element under the homomorphism $f_{\ast}:\pi_{1}(M,v^{\prime})\to\pi_{1}(S^{1},\bar{v})$. But instead this class is mapped by $f_{\ast}$ to the class of $(\bar{\alpha}_{1},\bar{\beta}_{1},\ldots,\bar{\alpha}_{n},\bar{\beta}_{n})$, a non-trivial element of $\pi_{1}(S^{1},\bar{v})$. ∎
We now prove Corollary 1.5.
Corollary 1.5.
Let $G\curvearrowright T$ be a non-trivial splitting with minimal action, with $G$ one-ended and finitely presented. Suppose the edge stabilizers are finitely generated, and suppose there is some edge stabilizer $G_{e}$ with more than one end. If the two halfspaces of $G\curvearrowright T$ associated to $e$ are one-ended then $H^{2}(G,\mathbb{Z}G)\neq\{0\}$.
Proof.
Following Definition 2, fix an orbit map $f:G\to VT$ and fix a finite generating set $S$ for $G$. Write $X$ for the Cayley graph of $G$ with respect to $S$. The halfspaces in $T$ bounded by the edge midpoint $\hat{e}$ have preimages in $G$ that induce halfspaces $\mathfrak{h}_{1}$ and $\mathfrak{h}_{2}$ in $X$. As discussed in Definition 2, $\mathfrak{h}_{1}$ and $\mathfrak{h}_{2}$ are connected subgraphs of $X$ for suitable choice of the finite generating set $S$. By hypothesis, $\mathfrak{h}_{1}$ and $\mathfrak{h}_{2}$ are one-ended.
The halfspaces $\mathfrak{h}_{1}$ and $\mathfrak{h}_{2}$ are disjoint, but contain all the vertices of $X$. Let $X_{1}$ and $X_{2}$ be the $L$-neighborhoods of $\mathfrak{h}_{1}$ and $\mathfrak{h}_{2}$ respectively, with $L\geq 1$ an integer. Then $X=X_{1}\cup X_{2}$, and $X_{1}$ and $X_{2}$ are one-ended (they are quasi-isometric to $\mathfrak{h}_{1}$ and $\mathfrak{h}_{2}$).
We now claim that $X_{1}\cap X_{2}$ is connected for large enough $L$, and that $G_{e}$ acts cocompactly on $X_{1}\cap X_{2}$. Let $\Sigma_{e}$ be the set of edges $(x_{1},x_{2})$ in $X$ with $x_{i}\in\mathfrak{h}_{i}$. For $(x_{1},x_{2})\in\Sigma_{e}$ we have $f(x_{1})$ and $f(x_{2})$ separated by $e$ in $T$. Let $m(x_{1},x_{2})$ denote the pair of integers $(d(f(x_{1}),e),d(f(x_{2}),e))$. There is a uniform bound on these integers, since there is a uniform bound on $d(f(x_{1}),f(x_{2}))$ for $(x_{1},x_{2})$ ranging over all edges in $X$. Moreover, if $g\in G$ and $(x_{1},x_{2}),(gx_{1},gx_{2})\in\Sigma_{e}$ with $m(x_{1},x_{2})=m(gx_{1},gx_{2})$, then $g\in G_{e}$. As a result, $\Sigma_{e}$ is a union of finitely many $G_{e}$-orbits of edges. Since $X$ is locally finite, it follows that $G_{e}$ acts cocompactly on $X_{1}\cap X_{2}$. As $G_{e}$ is finitely generated, it is also easy to see that $X_{1}\cap X_{2}$ is connected for large enough $L$.
Now add 2-cells to $X$ according to a finite presentation for $G$. This makes $X$ simply connected, and $G=\pi_{1}(X/G)$. Furthermore, for large enough $L$, each 2-cell will have attaching map contained entirely in either $X_{1}$ or $X_{2}$, so we may add each 2-cell to either $X_{1}$ or $X_{2}$ without increasing their 1-skeleta. After adding 2-cells, we still have $X=X_{1}\cup X_{2}$, and $X_{1}$ and $X_{2}$ are still one-ended.
By hypothesis, $G_{e}$ has more than one end, hence there is a finite subcomplex $K$ of $X$ such that $(X_{1}\cap X_{2})-K$ has more than one unbounded component. We apply Theorem 1.4 to deduce that $X$, and hence $G$, does not have pro-finite first homology at infinity. Finally, Theorem 5.3 implies that $H^{2}(G,\mathbb{Z}G)\neq\{0\}$. ∎
7 Artificial splittings and an example
We give an example of a finitely presented one-ended simply connected at infinity group that splits non-trivially over a finitely generated infinite-ended group. So there is no result that implies a general non-trivial splitting of a one-ended finitely presented group over an infinite-ended group is not simply connected at infinity. The next lemma describes how a splitting of a group can “artificially” induce a splitting over a larger subgroup.
Lemma 7.1.
Suppose $G$ splits as $A\ast_{C}B$ and $D$ is a subgroup of $B$ containing $C$. Then $G$ also splits as $(A\ast_{C}D)\ast_{D}B$.
Proof.
Let $\mathcal{P}_{Q}=\langle S_{Q}:R_{Q}\rangle$ be a presentation for $Q$, where $Q\in\{A,B,C,D\}$. Assume that $\mathcal{P}_{C}$ is a sub-presentation of $\mathcal{P}_{A}$ and $\mathcal{P}_{D}$ (so that $S_{C}\subset S_{A}$, $R_{C}\subset R_{A}$, $S_{C}\subset S_{D}$ and $R_{C}\subset R_{D}$). Also assume that $\mathcal{P}_{D}$ is a sub-presentation of $\mathcal{P}_{B}$. Then a presentation for $G$ is
$\mathcal{P}_{G}=\langle S_{A}\cup S_{B}:R_{A}\cup R_{B}\rangle.$
A presentation for $(A\ast_{C}D)\ast_{D}B$ is
$\langle S_{A}\cup S_{D}\cup S_{B}:R_{A}\cup R_{D}\cup R_{B}\rangle.$
This last presentation is $\mathcal{P}_{G}$ since $S_{D}\subset S_{B}$ and $R_{D}\subset R_{B}$. ∎
Example 7.2.
Let $\mathbb{Z}_{x}$ be the infinite cyclic group with generator $x$. Let $A=\mathbb{Z}_{a}\times\mathbb{Z}_{b}\times\mathbb{Z}_{f}=\mathbb{Z}^{3}$, $B=((\mathbb{Z}_{a}\times\mathbb{Z}_{b})\ast\mathbb{Z}_{c})\times\mathbb{Z}_{d}\times\mathbb{Z}_{e}$ and $C=\mathbb{Z}_{a}\times\mathbb{Z}_{b}$. The groups $A$ and $B$ are simply connected at infinity by [Jac82a, Theorem 2] and $C$ is one-ended. So $G=A\ast_{C}B$ is one-ended and simply connected at infinity by [Jac82b, Theorem 2].
Let $D=(\mathbb{Z}_{a}\times\mathbb{Z}_{b})\ast\mathbb{Z}_{c}$. The group $G$ artificially decomposes as
$G=(A\ast_{C}D)\ast_{D}B$
so the one-ended simply connected at infinity group $G$ splits non-trivially over the infinite-ended group $D=(\mathbb{Z}_{a}\times\mathbb{Z}_{b})\ast\mathbb{Z}_{c}$. Note that $A\ast_{C}D=A\ast\mathbb{Z}_{c}$ so we also have
$G=(A\ast\mathbb{Z}_{c})\ast_{D}B.$
Hence, in order to recognize this last splitting as an artificial splitting of $G$ we would have to recognize that $A\ast\mathbb{Z}_{c}$ splits non-trivially with $D$ as a vertex group. Observe that $A\ast_{C}D$ is an artificial splitting of $A\ast\mathbb{Z}_{c}$. Note also that the halfspace corresponding to $A*\mathbb{Z}_{c}$ in the last splitting has more than one end by Corollary 1.5.
8 Some examples of splittings with one-ended halfspaces
The following proposition provides a general construction of HNN extensions with one-ended halfspaces (without appealing to more advanced theory like JSJ splittings).
Proposition 8.1.
Suppose $G=A\ast_{C}$, $A$ and $C$ are finitely generated and infinite, $t$ is the stable letter and $t^{-1}ct=c$ for all $c\in C$. If $G$ is one-ended then all the halfspaces in this splitting are one-ended. In particular, if $G$ is also finitely presented and $C$ has more than one end, then $H^{2}(G,\mathbb{Z}G)\neq\{0\}$ (by Corollary 1.5).
Proof.
We use the construction of halfspaces from Definition 2. We have $S_{A}$ a finite generating set for $A$, and $S=S_{A}\cup\{t\}$ a finite generating set for $G$. The halfspace $\mathfrak{h}^{+}$ is the induced subgraph of $\Gamma=\operatorname{Cay}(G,S)$ with vertex set corresponding to elements of $G$ with normal forms that do not start with an element of $C$ followed by a negative power of $t$. The opposite halfspace is denoted $\mathfrak{h}^{-}$. If we choose $S$ so that it contains a generating set for $C$, then $\mathfrak{h}^{+}\cap\mathfrak{h}^{-}$ is connected and is isomorphic to a Cayley graph of $C$ (Remark 2.3). We denote $\mathfrak{h}^{+}\cap\mathfrak{h}^{-}$ by $\Gamma_{C}$. We show $\mathfrak{h}^{+}$ is one-ended (the argument for $\mathfrak{h}^{-}$ is similar). Suppose a finite subgraph $K$ of $\Gamma$ separates $\mathfrak{h}^{+}$, with at least two unbounded components $C_{1}$ and $C_{2}$. First observe that the intersection of $\Gamma_{C}$ with each of the components $C_{1}$ and $C_{2}$ is unbounded (otherwise $K$ union the intersection of this component with $\Gamma_{C}$ is a finite subgraph of $\Gamma$ separating $\Gamma$ with more than one unbounded component, but this implies $\Gamma$ has more than one end – contrary to our hypothesis). Now the product set $\{1,t,t^{2},\dots\}C\subset G$ spans a subgraph $\Delta$ of $\mathfrak{h}^{+}$, and $\Delta$ is a product of two infinite graphs (since $t$ commutes with every element of $C$), hence it is one-ended. But $C_{1}\cap\Delta$ and $C_{2}\cap\Delta$ are unbounded, hence they cannot be separated in $\Delta$ by the finite subgraph $K$, a contradiction. ∎
Next we show that a certain “doubling” process always leads to amalgamations with one-ended halfspaces whenever the double is one-ended. If $A$ is a group with subgroup $C$ then let $i:A\to A^{\prime}$ be an isomorphism such that $i(C)=C^{\prime}$. The double of $A$ across $C$ is the amalgamation $A\ast_{C=C^{\prime}}A^{\prime}$.
Proposition 8.2.
Suppose $A$ is a finitely generated group with finitely generated subgroup $C$. Let $G=A\ast_{C=C^{\prime}}A^{\prime}$ be the double of $A$ across $C$. If $G$ is one-ended, then the halfspaces of $G$ with respect to $C$ are one-ended. In particular, if $G$ is also finitely presented and $C$ has more than one end, then $H^{2}(G,\mathbb{Z}G)\neq\{0\}$ (by Corollary 1.5).
Proof.
First notice that there is a unique automorphism $r:G\to G$ satisfying $r(a)=i(a)$ for all $a\in A$ and $r(a^{\prime})=i^{-1}(a^{\prime})$ for all $a^{\prime}\in A^{\prime}$. If $X$ is the Cayley graph of $G$ with respect to a symmetric set of generators for $A$ and $A^{\prime}$ (and $C$ and $C^{\prime}$), then $r$ induces an involution $\hat{r}:X\to X$ that fixes the sub-Cayley graph for $C=C^{\prime}$ and exchanges halfspaces. More precisely, if $\mathfrak{h}=\mathfrak{h}_{A}$ and $\mathfrak{h}^{\prime}=\mathfrak{h}_{A^{\prime}}$ are the halfspaces of $X$ given by Definition 2, then $\hat{r}(\mathfrak{h})=\mathfrak{h}^{\prime}$ and $\hat{r}(\mathfrak{h}^{\prime})=\mathfrak{h}$. Suppose $K$ is a compact subset of $\mathfrak{h}$. Since $G$ (and hence $X$) is one-ended, there is a compact set $D\subset X$ such that any two points $x,y\in X-D$ can be joined by a path in $X-(K\cup\hat{r}(K))$. Suppose $x$ and $y$ are points in $\mathfrak{h}-D$ and $\alpha$ is a path in $X-(K\cup\hat{r}(K))$ from $x$ to $y$. Decompose $\alpha$ as $(\alpha_{1},\alpha_{1}^{\prime},\ldots,\alpha_{n},\alpha_{n}^{\prime})$ where $\alpha_{i}$ (respectively $\alpha_{i}^{\prime}$) has image in $\mathfrak{h}$ (respectively $\mathfrak{h}^{\prime}$). The path $(\alpha_{1},\hat{r}(\alpha_{1}^{\prime}),\ldots,\alpha_{n},\hat{r}(\alpha_{n}^{\prime}))$ is a path from $x$ to $y$ in $\mathfrak{h}-K$. Hence $\mathfrak{h}$ is one-ended (and $\mathfrak{h}^{\prime}$ similarly). ∎
The previous two propositions are good ways to produce examples of groups $G$ such that $H^{2}(G,\mathbb{Z}G)\neq\{0\}$, and hence groups that are not duality groups. We provide two such examples below (Examples 8.4 and 8.5). Recall that a group $G$ that acts freely and cocompactly on a contractible cell complex $X$ (by permuting cells) is a duality group of dimension $n$ if $H^{i}(G;\mathbb{Z}G)$ vanishes when $i\neq n$ and is torsion-free (as an abelian group) when $i=n$ (see R. Bieri and B. Eckmann’s article [BE73]).
Examples 8.4 and 8.5 are chosen in such a way that we cannot see a method other than using one-endedness of halfspaces to show that they are not duality groups. Before giving the examples, we discuss two other methods for showing that a group is not a duality group (which fail for our examples).
(Co)homological dimension: Any nontrivial amalgam $G=A*_{C}B$ or HNN extension $G=A*_{C}$ of an $n$-dimensional duality group $G$ requires $n-1\leq\mathrm{cd}(C)\leq\mathrm{cd}(A),\mathrm{cd}(B)\leq n$ [Bie81, Proposition 9.14]. Furthermore, if $G$ is a Poincaré duality group, then every subgroup of $G$ either has finite index or has homology dimension at most $n-1$ [Bie81, Proposition 9.22].
Mayer–Vietoris sequences: The Mayer–Vietoris long exact sequence can be applied to both amalgams and HNN extensions of groups. We state the result for amalgams.
Theorem 8.3 (Theorem 2.10, [Bie76]).
Suppose the group $G$ decomposes as $G=A\ast_{C}B$, where $A$ and $B$ are finitely presented and $C$ is finitely generated. Then one has a natural long exact sequence:
$\cdots\to H^{k}(G,\mathbb{Z}G)\to H^{k}(A,\mathbb{Z}G)\oplus H^{k}(B,\mathbb{Z}G)\to H^{k}(C,\mathbb{Z}G)\to H^{k+1}(G,\mathbb{Z}G)\to\cdots$
Suppose that the group $G$ decomposes as $G=A\ast_{C}B$, where $A$ and $B$ are one-ended and finitely presented and $C$ is finitely generated with more than one end. The number of ends of a finitely generated group is equal to $rank(H^{1}(G,RG))+1$ (see Theorem 13.5.5, [Geo08]). So we have $H^{1}(A,\mathbb{Z}G)=H^{1}(B,\mathbb{Z}G)=\{0\}$. Applying the above exact sequence, we obtain:
$0\to H^{1}(C,\mathbb{Z}G)\to H^{2}(G,\mathbb{Z}G)$
In particular, $H^{2}(G,\mathbb{Z}G)\neq\{0\}$ (which implies Jackson’s theorem). However, when $A$ or $B$ is not one-ended, this sequence gives little information about $H^{2}(G,\mathbb{Z}G)$.
Example 8.4.
Let $A=\mathbb{Z}^{n}$ ($n\geq 2$) with basis $\{a_{1},\ldots,a_{n}\}$. Consider the free product $A\ast\mathbb{Z}$, where the second factor is generated by an element $s$. Let $C$ be the subgroup of $A\ast\langle s\rangle$ with generating set $\{a_{1},\ldots,a_{n-1},s^{2},sa_{1}sa_{2}\cdots sa_{n}s\}$. Note that $C\cong\mathbb{Z}^{n-1}\ast F_{2}$. Define
$G=(A\ast\langle s\rangle)\ast_{C}$
with stable letter $t$ such that $t^{-1}ct=c$ for all $c\in C$.
Note that the base group $A\ast\langle s\rangle$ has cohomological dimension $n$ and the edge group $C$ has cohomological dimension $n-1$, so $G$ looks like a duality group from the perspective of cohomological dimension. The point of including the element $as_{1}sa_{2}\cdots sa_{n}s$ in $C$ is to prevent other splittings of $G$ over groups of lower cohomological dimension.
Now we show that $G$ is one-ended. Assume not. As $G$ is torsion free, it splits non-trivially as $U\ast V$ with $U$ and $V$ non-trivial torsion free groups. The subgroup generated by $C$ and the stable letter $t$ has the form $C\times\langle t\rangle$ and so is one-ended. Its intersection with the one-ended group $A$ is $\langle a_{1},\ldots,a_{n-1}\rangle$, which is infinite. This implies the group generated by $A$, $t$, $s^{2}$ and $as_{1}sa_{2}\cdots sa_{n}s$ is one-ended and so must be contained in a conjugate of $U$ or $V$ (say $U$). Killing the normal closure of $U$ in $G$ leaves $V$. But killing $U$ also kills $A$, $t$ and $s^{2}$ and so $V$ is at most $\mathbb{Z}_{2}$, a contradiction.
The edge group $C$ is infinite-ended and $G$ is finitely presented, so $H^{2}(G,\mathbb{Z}G)\neq\{0\}$ by Proposition 8.1. Hence $G$ is not a duality group.
Example 8.5.
Consider the group
$A=(F_{2}\times F_{2}\times F_{2})\ast F_{2}=(\langle x_{1},x_{2}\rangle\times\langle y_{1},y_{2}\rangle\times\langle z_{1},z_{2}\rangle)\ast\langle a_{1},a_{2}\rangle.$
Let $G=A\ast_{C=C^{\prime}}A^{\prime}$ be the double of $A$ across $C$, where
$C=\langle x_{1},y_{1},z_{1},a_{1}a_{2}a_{1}^{-1}a_{2}^{-1},a_{1}x_{2}a_{2}y_{2}a_{1}z_{2}\rangle.$
Note that $A$, $A^{\prime}$ and $C\cong\mathbb{Z}^{3}\ast F_{2}$ have cohomological dimension 3. Also note that $A$ and $A^{\prime}$ are infinite-ended, so we cannot use the Mayer–Vietoris sequence to show that $H^{2}(G,\mathbb{Z}G)\neq\{0\}$.
Now we show $G$ is one-ended. Suppose not. As $G$ is torsion free, it splits non-trivially as $U\ast V$. The subgroup of $G$ generated by
$\{x_{1}=x^{\prime}_{1},x_{2},x^{\prime}_{2},y_{1}=y^{\prime}_{1},y_{2},y^{\prime}_{2},z_{1}=z^{\prime}_{1},z_{2},z^{\prime}_{2}\}$
is isomorphic to $F_{2}^{3}*_{\mathbb{Z}^{3}}F_{2}^{3}$, so is one-ended (where $x^{\prime}_{i},y^{\prime}_{i},z^{\prime}_{i}$ are the images of $x_{i},y_{i},z_{i}$ under the isomorphism $A\to A^{\prime}$), as is the subgroup generated by
$\{a_{1},a_{2},a_{1}^{\prime},a_{2}^{\prime}\}$
(a surface group). These subgroups generate $G$. One lies in a conjugate $gUg^{-1}$, the other in a conjugate of $hVh^{-1}$ (killing the normal closure of $U$ leaves $V$). By a standard ping-pong argument in the Bass–Serre tree of the free product $U\ast V$, the subgroup $\langle gUg^{-1},hVh^{-1}\rangle$ splits as a free product $gUg^{-1}\ast hVh^{-1}$. But then the relator $a_{1}x_{2}a_{2}y_{2}a_{1}z_{2}(a_{1}^{\prime}x_{2}^{\prime}a_{2}^{\prime}y_{2}^{\prime}a_{1}^{\prime}z_{2}^{\prime})^{-1}$ has length 11 with respect to the normal form in $gUg^{-1}\ast hVh^{-1}$, which is impossible. Instead $G$ is one-ended.
The edge group $C$ is infinite-ended and $G$ is finitely presented, so Proposition 8.2 implies that the halfspaces of $G=A\ast_{C=C^{\prime}}A^{\prime}$ are one-ended, and that $H^{2}(G,\mathbb{Z}G)\neq\{0\}$. In particular $G$ is not a duality group.